### v1

# heightadjust=object, valign=c

Abstract
We study an old question in combinatorial group theory which can be traced back to a conjecture of Graham from 1971. Given a group $\Gamma$, and some subset $S\subseteq\Gamma$, is it possible to permute $S$ as $s_{1},s_{2},\ldots,s_{d}$ so that the partial products $\prod_{1\leq i\leq t}s_{i}$, $t\in[d]$ are all distinct? Most of the progress towards this problem has been in the case when $\Gamma$ is a cyclic group. We show that for any group $\Gamma$ and any $S\subseteq\Gamma$, there is a permutation of $S$ where all but a vanishing proportion of the partial products are distinct, thereby establishing the first asymptotic version of Graham’s conjecture under no restrictions on $\Gamma$ or $S$.
To do so, we explore a natural connection between Graham’s problem and the following very natural question attributed to Schrijver. Given a $d$-regular graph $G$ properly edge-coloured with $d$ colours, is it always possible to find a rainbow path with $d-1$ edges? We settle this question asymptotically by showing one can find a rainbow path of length $d-o(d)$. While this has immediate applications to Graham’s question for example when $\Gamma=\mathbb{F}_{2}^{k}$, our general result above requires a more involved result we obtain for the natural directed analogue of Schrijver’s question.
1 Introduction
In this paper, we study the following natural question in combinatorial number theory, first raised by Graham [29] in 1971, and reiterated by Erdős and Graham in 1980 [24].
Conjecture 1.1 (Graham, 1971).
For any $p$ prime and $a_{1},a_{2},\dots,a_{d}$ non-zero distinct elements of $\mathbb{Z}_{p}$, there exists a rearrangement of the elements as $a_{i_{1}},a_{i_{2}},\dots,a_{i_{d}}$ such that all partial sums $\sum_{j=1}^{t}{a_{i_{j}}}$, $1\leq t\leq d$ are distinct.
In his original paper, Graham [29] poses the rearrangement problem in conjunction with several other related problems concerning sumset structures, that is, the structure of the set of all numbers that can be formed by taking sums of elements from a given subset, which can be considered the principal domain of additive combinatorics. However, later papers of Graham [18, 16, 30] suggest that the rearrangement problem was at least partially motivated by practical applications to juggling.
The goal of the present paper is to provide asymptotic solutions to Graham’s conjecture and several other related problems. To achieve this, we primarily use the lens of extremal and probabilistic combinatorics; in particular, we exploit a connection between Graham’s conjecture and the rich area of finding rainbow subgraphs. An edge colouring of an undirected graph is proper if no two edges sharing a vertex have the same colour. In the directed setting, no pair of edges with a common start point and no pair of edges with a common end point may be monochromatic. In both cases, a subgraph of a coloured graph or digraph is rainbow if all its edges have distinct colours. Given a subset $S$ of a group $\Gamma$, we define the (coloured) Cayley graph on $\Gamma$ with generator set $S$ to be the edge-coloured directed graph $\mathrm{Cay}(\Gamma,S)$ on vertex set $\Gamma$ with an edge from $a$ to $ag$ of colour $g$ for every $a\in\Gamma$ and $g\in S$ [^1] Here and throughout the paper, we use additive notation if the underlying group is assumed to be abelian, and multiplicative notation otherwise.. If $S=\{a_{i_{1}},\ldots,a_{i_{d}}\}$ is a rearrangement of $S\subseteq\mathbb{Z}_{p}\setminus\{0\}$ with distinct partial sums, we see that $(a_{i_{1}},a_{i_{1}}+a_{i_{2}},\ldots,a_{i_{1}}+\cdots+a_{i_{d}})$ is a rainbow directed path in $\mathrm{Cay}(\mathbb{Z}_{p},S)$ with $d-1$ edges. Conversely, any rainbow directed path in $\mathrm{Cay}(\mathbb{Z}_{p},S)$ with $d-1$ edges [^2] Henceforth, in the context of directed graphs, a path always refers to a directed path, and in both directed and undirected settings, the length of a path refers to the number of edges it contains. is of the form $(x+a_{i_{1}},x+a_{i_{1}}+a_{i_{2}},\ldots,x+a_{i_{1}}+\cdots+a_{i_{d}})$ for some $x\in\mathbb{Z}_{p}$ and rearrangement $S=\{a_{i_{1}},\ldots,a_{i_{d}}\}$ with distinct partial sums. Before saying more about our methods, we give a survey of what is known about Graham’s conjecture as well as various generalisations thereof.
Most of the progress towards Graham’s conjecture has been in the cases when the generator set $S=\{a_{1},a_{2},\dots,a_{d}\}$ is very small or very large. Indeed, summarising the work of many, an approach based on the Combinatorial Nullstellensatz [1] verifies the conjecture when $|S|\leq 12$ and direct, constructive arguments can be used when $|S|\geq p-3$, see [20, 37] and references therein. It also follows from the results of [53] (see Section 6.2) that Graham’s conjecture is true for sets $S$ with $|S|=(1-o(1))p$. Recently, Kravitz [46] and independently Sawin [64] showed that Graham’s conjecture holds when $|S|\leq\log p/\log\log p$. In a subsequent paper, Bedert and Kravitz [11] significantly improved this bound by showing that the conjecture holds when $|S|\leq e^{(\log{p})^{1/4}}$, which has been an important milestone as this bound overcomes a natural barrier for the rectification techniques used in [46, 64].
Given an arbitrary group $\Gamma$ and some subset $S=\{a_{1},\ldots,a_{d}\}\subseteq\Gamma$, we say $S$ is rearrangeable if it is possible to order the elements as of $S$ as $a_{i_{1}},a_{i_{2}},\dots,a_{i_{d}}$ such that all partial products $\prod_{j=1}^{t}{a_{i_{j}}}$, $1\leq t\leq d$ are distinct. The observation mentioned earlier which connects rearrangeable sets in $\mathbb{Z}_{p}$ to rainbow paths in Cayley graphs extends to this more general context. That is, $S\subseteq\Gamma$ is rearrangeable if and only if $\mathrm{Cay}(\Gamma,S)$ contains a rainbow path of length $|S|-1$. Graham’s conjecture states that all subsets of $\mathbb{Z}_{p}\setminus\{0\}$ are rearrangeable, but in fact, the study of rearrangeable sets has a rich history that predates Graham’s conjecture. In 1961, Gordon [28] introduced the notion of sequenceable groups as those groups $\Gamma$ for which the set $S=\Gamma$ is rearrangeable. In the same paper, he characterised all sequenceable abelian groups. A conjecture by Keedwell [40] from 1981 says that the only non-abelian non-sequenceable groups are the dihedral groups of order $6$, $8$ and the quaternion group. This has recently been confirmed for all sufficiently large groups by Müyesser and Pokrovskiy [53]. In the same spirit, Ringel [58], motivated by some constructions arising in his celebrated proof of the Heawood map colouring conjecture [59], raised the question of which groups can be ordered as $a_{1},a_{2},\ldots,a_{n}$ where $a_{1}$ is the identity, the partial products $a_{1}$, $a_{1}a_{2}$, $\ldots$ , $a_{1}a_{2}\cdots a_{n-1}$ are all distinct, and $a_{1}a_{2}\cdots a_{n}=e$; that is, for which groups $\Gamma$ does $\mathrm{Cay}(\Gamma,\Gamma)$ contain a rainbow cycle of length $|\Gamma|-1$? This latter problem has similarly been resolved for large groups in [53]. The motivation for these kinds of problems comes from combinatorial design theory [26]. For example, a Latin square is an $n\times n$ array filled with $n$ symbols such that each symbol appears exactly once in each row and once in each column. A Latin square is called complete if every pair of distinct symbols appears exactly once in each order in adjacent horizontal cells and exactly once in adjacent vertical cells. It is not hard to see that any sequenceable group admits a Cayley (multiplication) table which is a complete Latin square. This connection ties sequenceability to decompositions of directed graphs into Hamiltonian paths [54], Heffter arrays [55] and even to experimental designs [10].
Showing rearrangeability of arbitrary subsets of any group turns out to be significantly more difficult, and here there has only been limited progress, even in the simplest case of $\Gamma=\mathbb{Z}_{p}$, as in the setting of Graham’s conjecture. We remark that a generalised version of Graham’s conjecture for all cyclic groups was posed by Archdeacon, Dinitz, Mattern, and Stinson [7]. A slightly stronger conjecture is due to Alspach [13, 21] who conjectured that if in Conjecture 1.1 we have an additional assumption that $\sum_{i=1}^{d}{a_{i}}\neq 0$, then all the partial sums are not only distinct but also non-zero. This is equivalent to saying that for any subset $S\subseteq\mathbb{Z}_{n}\setminus\{0\}$, $\mathrm{Cay}(\mathbb{Z}_{n},S)$ contains either a rainbow path of length $|S|$ or a rainbow cycle of length $|S|$. Alspach was motivated by applications to cycle decompositions of complete graphs [3], complete graphs plus or minus a $1$-factor [62, 63] and complete symmetric digraphs [4]. Various versions of Alspach’s conjecture have been reiterated and stated, for example for general abelian groups by Costa, Morini, Pasotti and Pellegrini [22], and by Costa, Fiore, Ollis and Rovner-Frydman [21]. Below we state a version of Graham’s conjecture for general abelian groups which is due to Alspach and Liversidge [5].
Conjecture 1.2.
For any finite abelian group $\Gamma$, any subset $S\subseteq\Gamma\setminus\{0\}$ is rearrangeable.
Conjecture 1.2 is known to hold when $|S|\leq 9$ [5]; interestingly, the proof uses a method based on posets which is distinct from the other methods mentioned above for the $\Gamma=\mathbb{Z}_{p}$ case. Our first result gives an approximate answer to these conjectures, and moreover, it applies to non-abelian groups as well.
Theorem 1.3.
For any finite group $\Gamma$ and any subset $S\subseteq\Gamma$ there exists an ordering of elements of $S$ in which at least $(1-o(1))|S|$ many partial products are distinct.
Theorem 1.3 guarantees that all subsets are approximately rearrangeable in a very general setup. However, there is yet another natural notion of an approximate rearrangeability introduced by Archdeacon, Dinitz, Mattern, and Stinson (see Problem 1 in [7]), where given $S$, we search for a subset $S^{\prime}\subseteq S$ as large as possible such that $S^{\prime}$ is rearrangeable. A greedy argument easily produces such an $S^{\prime}$ of size $|S|/2$ (see the work of Hicks, Ollis, and Schmitt [37] for slightly better bounds). However, a major milestone here would be to find a rearrangeable subset $S^{\prime}$ of size $(1-o(1))|S|$. Observe that the existence of such $S^{\prime}$ implies that $S$ can be rearranged approximately in the sense of Theorem 1.3, simply by arbitrarily permuting $S\setminus S^{\prime}$. Moreover, it is feasible that such a strong approximate result could pave the way for exact results. Indeed, the aforementioned result of Müyesser and Pokrovskiy [53] showing that large non-abelian groups are sequenceable relies on a strong approximation which holds in the regime when $\Gamma$ and $S$ have comparable size. More broadly, Keevash’s [41] celebrated proof of the existence of designs, the recent proof of Ringel’s conjecture by Montgomery, Pokrovskiy, and Sudakov [49] (see also the independent proof of Keevash and Staden [44]), and the resolution of the Ryser-Brualdi-Stein conjecture by Montgomery [48] all rely on the approximate versions of the corresponding results  [60, 51, 43] and use a similar framework, commonly referred to as the absorption method, to obtain exact results from approximate ones.
We are able to achieve this strong approximation in three distinct settings, as summarised below.
Theorem 1.4.
For any group $\Gamma$ and any subset $S\subseteq\Gamma$ of size $d:=|S|$, there exists a rearrangeable subset $S^{\prime}\subseteq S$ of size $(1-o(1))d$ if any one of the following holds.
  1. (a)
    $S$ contains only involutions, that is, elements of order two.
  2. (b)
    $d=\Omega(|\Gamma|)$.
  3. (c)
    $d\geq p^{3/4+o(1)}$ and $\Gamma=\mathbb{Z}_{p}$ for some prime $p$.
Note that Theorem 1.4(a) implies that the same conclusion holds for any subset of the group $\mathbb{F}_{2}^{k}$. To prove Theorem 1.3 and Theorem 1.4, we use various techniques from the toolkit of probabilistic combinatorics, such as robust expansion, and in the case of Theorem 1.4(c), some tools from additive combinatorics, such as Pollard’s inequality, which is a strengthening of the classical Cauchy-Davenport inequality. Theorem 1.4(a) and Theorem 1.4(b) are of special interest, as they are both consequences of much more general results concerning the existence of rainbow paths in properly edge-coloured graphs and digraphs. Questions of this nature are both natural and have been the subject of considerable research, going all the way back to Euler’s work in 18th century on the existence of transversals in Latin squares. In the last decade alone, there have been several breakthroughs in this area. Most notably, the famous Ryser-Brualdi-Stein conjecture [61, 15, 66] from 1967 states that in any Latin square of order $n$ there exists a transversal of order $n-1$, that is $n-1$ cells which share no column, row or a symbol. This has equivalent formulations in terms of rainbow structures in graphs and digraphs. In the undirected setting, a Latin square of order $n$ corresponds to a proper edge colouring of the complete bipartite graph $K_{n,n}$ with $n$ colours, and a transversal corresponds to a rainbow matching. Last year, following a recent improvement on the best lower bound on the size of a rainbow matching of the form $n-o(n)$ by Keevash, Pokrovskiy, Sudakov and Yepremyan [43], Montgomery [48] resolved the conjecture in his tour de force work. In the directed setting, a Latin square of order $n$ corresponds to a proper edge colouring with $n$ colours of the complete symmetric digraph on $n$ vertices with loops at each vertex, and a transversal corresponds to a rainbow subgraph whose components are directed paths and cycles (see [12] for a more detailed explanation). There are also several related problems which inquire about the existence of rainbow subgraphs with a more specific structure; for example, subgraphs which are large path forests or long cycles (see [33, 12]) or cycle factors of a particular cycle type (see [27, 52]). We refer the reader to comprehensive surveys by Pokrovskiy [56], by Montgomery [50], and by Sudakov [67] for a broader overview of the area. For the rest of this paper, we will only focus on the existence of rainbow paths in properly edge-coloured graphs and digraphs.
In this direction, Hahn [34] conjectured in 1980 that in any proper edge colouring of $K_{n}$ there exists a rainbow path of length $n-1$. This was refuted by Maamoun and Meyniel [47], by considering $\mathrm{Cay}(\mathbb{F}_{2}^{k},\mathbb{F}_{2}^{k}\setminus\{0\})$ which has no rainbow Hamilton path. Andersen’s conjecture [6] from 1989 proposes a weakening of Hahn’s conjecture and states that there exists a rainbow path of length $n-2$ in any properly coloured $K_{n}$. After a lot of partial progress, Andersen’s conjecture has been proven asymptotically by Alon, Pokrovskiy, and Sudakov [2] who showed the existence of rainbow paths of length $n-o(n)$ (see the work of Balogh and Molla [9] for the current best lower bound).
Schrijver [65] asked for a far reaching generalisation of Andersen’s conjecture by postulating the existence of a rainbow path of length $d-1$ in any properly $d$-edge-coloured $d$-regular graph $G$, and he verified this conjecture whenever $d\leq 10$. The best general bound on Schrijver’s problem guarantees a path of length roughly $32d/$ 3, due to Chen and Li (unpublished, see [8]), and independently, Johnston, Palmer, and Sarkar [38]. Babu, Chandran, and Rajendraprasad [8] showed that if the graph is $C_{4}$-free then a rainbow path of length $d-o(d)$ exists, and moreover if the girth is of order $\Omega(\log{d})$, then a rainbow path of length $d-2$ exists. In work independent from ours, Conlon and Haenni [19] have asymptotically resolved Schrijver’s problem for random $d$-regular graphs. We resolve Schrijver’s problem asymptotically in full generality, by showing the following.
Theorem 1.5.
Any properly edge-coloured $d$-regular graph contains a rainbow path of length $(1-o(1))d$.
Observe that the $d=n-1$ case of Theorem 1.5 corresponds to Andersen’s conjecture. Thus, our result recovers the aforementioned asymptotic result of Alon, Pokrovskiy, and Sudakov [2] for complete graphs and extends it into a much more difficult setting where the host graph can be extremely sparse. Furthermore, Theorem 1.5 also directly implies Theorem 1.4(a), thereby giving a strong approximation for Conjecture 1.2 in the case of $\Gamma=\mathbb{F}_{2}^{k}$.
Given the connection between rearrangements and rainbow directed paths described above, it is natural to wonder if the algebraic structure of a coloured Cayley graph is relevant for Graham’s conjecture, or more generally, for Conjecture 1.2. Specifically, we ask the following, which may be considered to be a directed generalisation of Schrijver’s problem to directed $d$-regular digraphs, that is, digraphs in which every vertex has in-degree and out-degree exactly $d$.
Problem 1.5.
Let $G$ be a $d$-regular digraph properly edge-coloured with $d$ colours. Does $G$ contain a directed rainbow path with $d-1$ edges?
This is known to be true asymptotically for complete symmetric digraphs by results of Benzing, Pokrovskiy and Sudakov [12], mirroring the asymptotic results on Andersen’s conjecture [2]. An affirmative answer to Problem 1 would be quite consequential, as this would resolve Graham’s conjecture, Conjecture 1.2, and the natural generalisation of Conjecture 1.2 to non-abelian groups as well, and therefore even an approximate answer is of great interest. We are able to give such an approximate answer to Problem 1 in the dense regime.
Theorem 1.6.
Let $G$ be a $d$-regular, properly edge-coloured digraph on $n$ vertices, with $d=\Omega(n)$. Then, $G$ contains a rainbow path with $(1-o(1))d$ edges.
Note that the above implies Theorem 1.4(b) directly. A common theme in the proofs of Theorem 1.3 and Theorem 1.5 is a reduction of the general case to the dense regime, hence Theorem 1.6 plays a key role in our arguments. Of course, some graphs do not contain any dense spots and this translates to a strong expansion property that allows us to build a long rainbow path by an elementary argument, see Lemma 4.1 for illustration. The challenge, then, lies in graphs that are neither dense, nor strongly expanding. This points to a recurring difficulty in our arguments, namely, we have to work with mild forms of expansion whilst building long rainbow paths. To see an example of how we achieve this, we refer the reader to Lemma 6.3 which introduces a novel procedure that outputs a rainbow path of asymptotically the optimal length in graphs that satisfy a rather delicate notion of expansion.
Organisation. We introduce some notation and state a few standard results and observations in Section 2. Section 3 proves some versions of our main results where paths are replaced with path forests with few components. Section 4 introduces some notions of expansion and builds a toolkit for working with expanders. Section 5 proves Theorem 1.6 and Theorem 1.4(c) using the tools derived in Sections 3 and 4. Section 6 contains a proof of Theorem 1.5. Section 7 contains a proof of Theorem 1.3. In Section 8 we make some concluding remarks and suggest some directions for future research.
2 Notation and preliminaries
Notation. The digraphs we consider are loopless, and for each pair $(u,v)$ of distinct vertices, we allow at most one edge from $u$ to $v$, which we denote by $(u,v)$. We do, however, allow both edges $(u,v)$ and $(v,u)$ to appear in the same digraph. If $G$ is a (possibly edge-coloured) digraph, then for $U,V\subseteq V(G)$, we write $e_{G}(U,V)$ to denote the number of edges $(u,v)$ with $u\in U$ and $v\in V$. As special cases, for a vertex $v\in V(G)$, we denote the out-degree of $v$ by $\deg^{+}_{G}(v):=e_{G}(\{v\},V(G))$ and the in-degree of $v$ by $\deg^{-}_{G}(v):=e_{G}(V(G),\{v\})$. We also write $\deg_{G}^{+}(v,U):=e_{G}(\{v\},U)$, $\deg_{G}^{-}(v,U):=e_{G}(U,\{v\})$, $\partial^{+}_{G}(U):=e_{G}(U,V(G)\setminus U)$, and $\partial^{-}_{G}(U):=e_{G}(V(G)\setminus U,U)$. We omit the subscript $G$ when the digraph $G$ is clear from context.
[…]
A digraph is called symmetric if every edge is contained in a $2$-cycle. To an undirected simple graph $G$, we associate a symmetric digraph $\hat{G}$ by replacing each undirected edge $uv\in E(G)$ with the two directed edges $(u,v)$ and $(v,u)$. If $G$ is edge-coloured, then both $(u,v)$ and $(v,u)$ are given the colour of $uv$. The resulting edge colouring of $\hat{G}$ is proper if and only if the edge colouring of $G$ is proper. Further, the rainbow paths (and walks) in $\hat{G}$ are in one-to-one correspondence with the rainbow paths (and walks) in $G$. Thus, for our purposes, we can treat $G$ and $\hat{G}$ as the same object when it is convenient; that is, we consider undirected graphs as special cases of directed graphs. We use the same notation for undirected graphs as for directed graphs described above, except that we omit the superscript ‘$+$’ or ‘--’ (for example, for a vertex $v\in V(G)$, we write $\deg_{G}(v):=\deg^{+}_{\hat{G}}(v)=\deg^{-}_{\hat{G}}(v)$ for the degree of $v$).
We use “$\alpha\ll\beta$ implies $P(\alpha,\beta)$” as shorthand to denote that there exists an increasing function $f$ such that for any $\beta$, $P(\alpha,\beta)$ holds for $\alpha\leq f(\beta)$.
Lemma 2.1 (Chernoff’s inequality).
Let $X$ be a sum of independent Bernoulli random variables with $\mathbb{E}(X)=\mu$. Then for every $t>0$,
  • $\mathbb{P}(X\leq\mu-t)\leq\exp(-t^{2}/(2\mu))$;
  • $\mathbb{P}(X\geq\mu+t)\leq\exp(-t^{2}/(2\mu+t))$.
We conclude the section with two simple observations, which extend classical arguments from the non-rainbow setting for finding long paths.
Observation 2.2.
A properly edge-coloured digraph with minimum out-degree $d$ contains a rainbow path of length at least $d/2$.
  • Proof.
    Let $P$ be a rainbow path of maximum length $\ell$, and let $v$ be the terminal vertex of $P$. Let $C^{\prime}$ be the set of colours not in $C(P)$. By maximality of $P$, every $C^{\prime}$-coloured out-edge from $v$ must go to $V(P)\setminus\{v\}$, so $d-\ell\leq\deg(v;C^{\prime})\leq\ell$. Thus $\ell\geq d/2$. ∎
Observation 2.3.
A properly edge-coloured undirected graph with average degree $d$ contains a rainbow path of length at least $d/4$.
  • Proof.
    Iteratively delete all vertices with degree less than $d/2$ as long as possible. The number of edges deleted is less than $dn/2$, so the remaining graph is non-empty and by construction has minimum degree at least $d/2$. Now applying Observation 2.2 to the associated symmetric digraph gives a rainbow path of length at least $d/4$. ∎
3 Rainbow path forests with few components
A path forest in a (directed) graph $G$ is a subgraph of $G$ whose components are paths. In this section, we prove two auxiliary lemmas on the existence of large rainbow path forests with few components in properly edge-coloured graphs and digraphs. The first one only applies to undirected graphs and requires the maximum degree to be sublinear in the number of vertices, but provides a rainbow forest with number of edges asymptotic in average degree. The second lemma we applies in both the directed and undirected case. Further, it makes essentially no requirement on the degree, but it only provides a rainbow forest with number of edges asymptotic in minimum degree. This distinction will be important when passing to subgraphs where we might lose control of the minimum degree but keep control of the average degree.
The following is the first of our rainbow path forest lemmas, which is proved by an iterative application of Observation 2.3 in the undirected case.
Lemma 3.1.
For any $\varepsilon>0$ there exists $d_{0}$ such that for all $d\geq d_{0}$ the following holds. If $G$ is a properly edge-coloured graph on $n$ vertices with average degree $d$ and maximum degree $\Delta\leq\varepsilon n/4$, then $G$ contains a rainbow path forest with at most $\left\lceil\log_{8/7}{\varepsilon^{-1}}\right\rceil$ many components and at least $(1-\varepsilon)d$ edges.
  • Proof.
    Let $0\leq m\leq\left\lceil\log_{8/7}\varepsilon^{-1}\right\rceil$ be maximum such that there exists a rainbow path forest $\mathcal{P}$ with $m$ components and at least $(1-(7/8)^{m})d$ edges. Clearly the empty graph satisfies these conditions for $m=0$, so this is well-defined. It suffices to show that $|E(\mathcal{P})|\geq(1-\varepsilon)d$.
    Suppose not. Then $(1-(7/8)^{m})d\leq e(\mathcal{P})<(1-\varepsilon)d$, so $m\leq\lceil\log_{8/7}\varepsilon^{-1}\rceil-1$. Note that
    $|V(\mathcal{P})|=e(\mathcal{P})+m\leq(1-\varepsilon)d+\log_{8/7}\varepsilon^{-1}\leq d$
    since $1/d\ll\varepsilon$. Let $G^{\prime}$ be obtained from $G$ by removing the vertices and colours of $\mathcal{P}$, and let $d^{\prime}:=d-e(\mathcal{P})$, which satisfies $\varepsilon d\leq d^{\prime}\leq(7/8)^{m}d$ by assumption. Since $G$ is properly edge-coloured, $G\setminus C(\mathcal{P})$ has at least $nd^{\prime}/2$ edges, and at most $\Delta|V(\mathcal{P})|$ of these are incident to some vertex in $V(\mathcal{P})$. This leaves at least
    $nd^{\prime}/2-\Delta|V(\mathcal{P})|\geq nd^{\prime}/2-\varepsilon nd/4\geq nd^{\prime}/4$
    edges in $G^{\prime}$ since $\Delta\leq\varepsilon n/4$ and $d^{\prime}\geq\varepsilon d$. Consequently, $G^{\prime}$ has average degree at least $d^{\prime}/2$, so by Observation 2.3, $G^{\prime}$ has a rainbow path $P$ of length at least $d^{\prime}/8$. Now $\mathcal{P}\cup P$ is a rainbow path forest with $m+1\leq\left\lceil\log_{8/7}\varepsilon^{-1}\right\rceil$ components and at least
    $e(\mathcal{P})+\frac{d^{\prime}}{8}=d-\frac{7d^{\prime}}{8}\geq\left(1-\left(\frac{7}{8}\right)^{m+1}\right)d$
    edges since $d^{\prime}\leq(7/8)^{m}d$. This contradicts the maximality of $m$. The proof is complete. ∎
We proceed to our second rainbow path forest lemma. The proof of this lemma is motivated by ideas from [12].
Lemma 3.2.
Let $0<\varepsilon\leq 1$, and let $G$ be a properly edge-coloured digraph with minimum in-degree $\delta^{-}(G)\geq 9\varepsilon^{-3}$. Then $G$ contains a rainbow path forest with at most $9\varepsilon^{-2}$ components and at least $(1-\varepsilon)\delta^{-}(G)$ edges.
  • Proof.
    Set $q,s:=\lceil 2\varepsilon^{-1}\rceil\leq 3\varepsilon^{-1}$, and set $d:=\delta^{-}(G)$. Let $\mathcal{P}$ be a rainbow path forest with at most $qs$ components, with the maximum possible number of edges. Assume for the sake of contradiction that $e(\mathcal{P})<(1-\varepsilon)d$. Then since $qs\leq\varepsilon d$, we have
    $|V(\mathcal{P})|<(1-\varepsilon)d+qs\leq d\leq|V(G)|,$
    so by adding isolated vertices to $\mathcal{P}$ one by one if necessary, we may assume that $\mathcal{P}$ has exactly $qs$ components.
    Let $U$ be the set of vertices in $\mathcal{P}$ with in-degree 00, and let $W$ be the set of vertices in $\mathcal{P}$ with out-degree 00. For $v\in V(\mathcal{P})\setminus W$, let $C(v)$ denote the colour of the edge coming out of $v$ in $\mathcal{P}$.
    Let $\{Q_{0},\ldots,Q_{s-1}\}$ be a partition of $U$ into sets of size exactly $q$. Let $C_{0}:=C(G)\setminus C(\mathcal{P})$ denote the set of colours in $G$ that do not appear in $\mathcal{P}$. For $1\leq i\leq s$, define
    $V_{i}$ :=$:=\{v\in V(G):\deg^{+}(v,Q_{i-1};C_{i-1})\geq 2\};$
    $V_{i}^{\prime}$ :=$:=\{v\in V(G):\deg^{+}(v,Q_{i-1};C_{i-1})=1\};$
    $C_{i}$ $:=C_{i-1}\cup\{C(v):v\in V_{i}\cap V(\mathcal{P})\setminus W\}.$
    Claim.
    For every $1\leq i\leq s$, we have $V_{i}\subseteq V(\mathcal{P})\setminus W$ and $V_{i}^{\prime}\subseteq V(\mathcal{P})$.
    • Proof.
      Suppose not. Then there exists a vertex $v$ such that either $v\in W\cap V_{i}$ or $v\in(V_{i}\cup V_{i}^{\prime})\setminus V(\mathcal{P})$. Let $1\leq j\leq i$ be minimum such that there exists a rainbow path forest $\mathcal{P}^{\prime}$ satisfying the following properties:
      1. (i)
        the number of components in $\mathcal{P}^{\prime}$ is at most $qs+1$;
      2. (ii)
        $C(\mathcal{P}^{\prime})=C(\mathcal{P})$;
      3. (iii)
        every vertex in $Q_{0}\cup\cdots\cup Q_{j-1}$ has in-degree 00 in $\mathcal{P}^{\prime}$;
      4. (iv)
        every edge in $\mathcal{P}$ whose colour is in $C_{j-1}\setminus C_{0}$ is also an edge in $\mathcal{P}^{\prime}$;
      5. (v)
        there exists a vertex $v^{\prime}\in V(\mathcal{P}^{\prime})$ with out-degree 00 in $\mathcal{P}^{\prime}$ such that either $v^{\prime}\in V_{j}$, or $v^{\prime}\in V_{j}^{\prime}$ is isolated in $\mathcal{P}^{\prime}$.
      Note that $\mathcal{P}\cup\{v\}$ satisfies these conditions for $j=i$ with $v$ playing the role of $v^{\prime}$, so $\mathcal{P}^{\prime}$ is well-defined.
      By definition, we have $v^{\prime}\in V_{j}\cup V_{j}^{\prime}$, so there exists an out-neighbour $u\in Q_{j-1}$ of $v^{\prime}$ such that $C(v^{\prime},u)\in C_{j-1}$. In fact, by (v), there are always at least two choices for such a $u$ unless $v^{\prime}$ is isolated in $\mathcal{P}^{\prime}$, so we can always ensure that $u$ and $v^{\prime}$ are in different components of $\mathcal{P}^{\prime}$.
      If $C(v^{\prime},u)\in C_{0}$, then $C(v^{\prime},u)\notin C(\mathcal{P})=C(\mathcal{P}^{\prime})$ by (ii), so adding the edge $(v^{\prime},u)$ to $\mathcal{P}^{\prime}$ produces a rainbow path forest with at most $qs$ components (by (i)) and with $|C(\mathcal{P})|+1$ edges, contradicting the maximality of $\mathcal{P}$.
      Thus, we may assume $C(v^{\prime},u)\in C_{j-1}\setminus C_{0}$, so by (iv), there exists an edge $(v^{\prime\prime},w)\in E(\mathcal{P})\cap E(\mathcal{P}^{\prime})$ with $C(v^{\prime\prime},w)=C(v^{\prime},u)$. Since $C(v^{\prime\prime},w)\in C_{j-1}\setminus C_{0}$, there exists $1\leq j^{\prime}<j$ such that $C(v^{\prime\prime})=C(v^{\prime\prime},w)\in C_{j^{\prime}}\setminus C_{j^{\prime}-1}$. By definition of $C_{j^{\prime}}$, this means that $v^{\prime\prime}\in V_{j^{\prime}}$. Now let $\mathcal{P}^{\prime\prime}$ be the rainbow path forest obtained from $\mathcal{P}^{\prime}$ by deleting the edge $(v^{\prime\prime},w)$ and adding the edge $(v^{\prime},u)$. Based on what we have observed, it is easily seen that $\mathcal{P}^{\prime\prime}$ satisfies properties (i)–(v) with $j^{\prime}$ playing the role of $j$ and $v^{\prime\prime}$ playing the role of $v^{\prime}$. This contradicts the minimality of $j$, and proves the claim. ∎
    Note that the claim implies for each $1\leq i\leq s$ that
    $|C_{i}\setminus C_{0}|\geq|V_{i}|,$ (1)
    and also
    $|V_{i}^{\prime}|\leq|V(\mathcal{P})|=e(\mathcal{P})+qs\leq d.$ (2)
    We now prove a lower bound on $|V_{i}|$ in terms of $|C_{i-1}\setminus C_{0}|$ by double counting the quantity $e(V(G),Q_{i-1};C_{i-1})$. First, because $C(G)\setminus C(\mathcal{P})=C_{0}\subseteq C_{i-1}$, the number of colours in $C(G)\setminus C_{i-1}$ is only $|C(\mathcal{P})|-|C_{i-1}\setminus C_{0}|\leq(1-\varepsilon)d-|C_{i-1}\setminus C_{0}|$. Therefore, since $G$ is properly edge-coloured with minimum in-degree $d$, we have $\deg^{-}(u;C_{i-1})\geq|C_{i-1}\setminus C_{0}|+\varepsilon d$ for every $u\in Q_{i-1}$, hence
    $e(V(G),Q_{i-1};C_{i-1})\geq(|C_{i-1}\setminus C_{0}|+\varepsilon d)q.$
    On the other hand, by (2) and our choice of $q$, we have
    $(|C_{i-1}\setminus C_{0}|+\varepsilon d)q$ $\leq e(V(G),Q_{i-1};C_{i-1})$
    $=e(V_{i},Q_{i-1};C_{i-1})+e(V_{i}^{\prime},Q_{i-1};C_{i-1})$
    $\leq|V_{i}|q+|V_{i}^{\prime}|$
    $\leq(|V_{i}|+\varepsilon d/2)q.$
    Rearranging gives
    $|V_{i}|\geq|C_{i-1}\setminus C_{0}|+\varepsilon d/2.$
    Combining this with (1) and applying induction, we have
    $|C_{s}\setminus C_{0}|\geq s\varepsilon d/2\geq d$
    since $s\geq 2\varepsilon^{-1}$. But $C_{s}\setminus C_{0}\subseteq C(\mathcal{P})$, so it can only have size at most $(1-\varepsilon)d$, a contradiction. ∎
4 Expansion-based tools
In this section, we establish a number of useful properties of a certain type of expander graphs. We start by proving some introductory results for much stronger notions of expansion in Section 4.1, in part as a warm-up to introduce easier variants of some of our new ideas. In Section 4.2, we introduce the notion of robust expansion. We show how to find a robust expander in a general nearly regular digraph and establish the key property of robust expanders that we will use.
4.1 Long rainbow paths in strong expanders
In this section, we show how to find long rainbow paths in digraphs with very strong expansion properties. The main lemma of this section (Lemma 4.1) is not used later in our paper, but we present it as a warm-up to highlight the connection between expansion and long rainbow paths. In particular, this lemma implies that an $n$-vertex random $d$-regular (directed) graph with $d=o(n)$ contains a rainbow path of length $d-o(d)$ with high probability. We learned that Conlon and Haenni [19] independently proved a similar lemma for this purpose in the undirected setting, using a Pósa rotation argument. Our proof is slightly different, and it applies in both directed and undirected settings. Later on, we prove another lemma (Lemma 5.1) which implies the same conclusion for dense random $d$-regular (directed) graphs.
Definition.
A digraph $G$ is $(\varepsilon,d)$-locally sparse if every set $U\subseteq V(G)$ of size at most $d$ satisfies $e_{G}(U,U)\leq\varepsilon d^{2}$.
Lemma 4.1.
Let $\varepsilon\in(0,1)$, and let $d\in\mathbb{N}$. Let $G$ be a properly edge-coloured digraph of minimum out-degree $d$. If $G$ is $(\varepsilon^{2}/4,d)$-locally sparse, then $G$ has a rainbow directed path of length at least $(1-\varepsilon)d$.
  • Proof.
    Define an $(\ell,k)$-broom in $G$ to be a rainbow subgraph $H=P\cup T$ of $G$, where $P$ is a path of length $\ell$ ending at some vertex $v$, and $T$ is an outward-oriented star with $k$ edges centred at $v$, with $V(P)\cap V(T)=\{v\}$. Clearly $G$ has a $(0,\lceil\varepsilon d/2\rceil)$-broom since $\delta^{+}(G)\geq d\geq\lceil\varepsilon d/2\rceil$.
    Consider an $(\ell,\lceil\varepsilon d/2\rceil)$-broom $H=P\cup T$ in $G$ with $\ell$ as large as possible. Suppose, towards a contradiction, that $G$ has no rainbow path of length at least $(1-\varepsilon)d$, so $\ell<(1-\varepsilon)d-1$. Let $v$ be the centre of $T$. For each vertex $u\in V(T)\setminus\{v\}$, there are at least $\varepsilon d$ colours leaving $u$ which do not appear among $E(P)\cup\{(v,u)\}$. Let $S_{u}$ be this set of colours, and let
    $N_{u}:=\{w\in N^{+}(u):C(u,w)\in S_{u}\}$
    be the out-neighbourhood of $u$ in these colours. By the maximality of $H$, we have
    $|N_{u}\setminus V(P)|<\varepsilon d/2,$
    for every $u\in V(T)\setminus\{v\}$, or else we would have an $(\ell+1,\lceil\varepsilon d/2\rceil)$-broom. This gives
    $e(V(H),V(H))\geq e(V(T)\setminus\{v\},V(P))\geq\sum_{u\in V(T)\setminus\{v\}}|N_{u}\cap V(P)|\geq\varepsilon^{2}d/4.$
    But $|V(H)|<(1-\varepsilon)d-1+\lceil\varepsilon d/2\rceil\leq d$, so $G$ fails to be $(\varepsilon^{2}/4,d)$-locally sparse, a contradiction. ∎
4.2 Robust expanders, how to find them, and how to use them
In this section, we build a toolkit for finding long rainbow paths in dense graphs and digraphs by way of the following robust expansion property.
Definition.
Let $G$ be a directed graph on $n$ vertices. For $U\subseteq V(G)$ and $\nu>0$, we define the $\nu$-robust out-neighbourhood of $U$ in $G$ to be the set
$RN_{\nu,G}^{+}(U):=\{v\in V(G):|N^{-}(v)\cap U|\geq\nu n\}.$
We say that $G$ is a robust $(\nu,\tau)$-out-expander if every $U\subseteq V(G)$ with $\tau n\leq|U|\leq(1-\tau)n$ satisfies
$|RN_{\nu,G}^{+}(U)\setminus U|\geq\nu n.$
Similarly, we define the $\nu$-robust in-neighbourhood of $U$ in $G$ to be the set
$RN_{\nu,G}^{-}(U):=\{v\in V(G):|N^{+}(v)\cap U|\geq\nu n\}.$
We say that $G$ is a robust $(\nu,\tau)$-in-expander if every $U\subseteq V(G)$ with $\tau n\leq|U|\leq(1-\tau)n$ satisfies
$|RN_{\nu,G}^{-}(U)\setminus U|\geq\nu n.$
The following lemma shows how to find such a robust expander subgraph with essentially the same degrees in any digraph with linear minimum degree and similar in- and out-degrees at every vertex.
Lemma 4.2.
For any $\delta,\tau,\alpha>0$, there exist $\gamma,\nu>0$ such that the following holds for every positive integer $n$ and every $n$-vertex digraph $G$. Suppose that $\delta^{\pm}(G)\geq\alpha n$ and that $|\deg^{+}(v)-\deg^{-}(v)|\leq\gamma n$ for every $v\in V(G)$. Let $w:V(G)\to\mathbb{R}_{+}$ be any weight function on the vertices of $G$. Then there exists a nonempty induced subgraph $G^{\prime}$ of $G$ satisfying the following:
  1. A1
    $G^{\prime}$ is a robust $(\nu,\tau)$-out-expander and a robust $(\nu,\tau)$-in-expander;
  2. A2
    every vertex $v\in V(G^{\prime})$ satisfies $\deg_{G^{\prime}}^{+}(v)\geq\deg_{G}^{+}(v)-\delta n$ and $\deg_{G^{\prime}}^{-}(v)\geq\deg_{G}^{-}(v)-\delta n$.
  3. A3
    $\frac{1}{|V(G^{\prime})|}\sum_{v\in V(G^{\prime})}w(v)\leq 2\frac{1}{n}\sum_{v\in V(G)}w(v)$.
Remark. The preceding lemma may be applied to undirected graphs with minimum degree $\Omega(n)$ as well, by way of their associated symmetric digraphs. The condition that $|\deg^{+}(v)-\deg^{-}(v)|\leq\gamma n$ is satisfied automatically for any $\gamma$. For this special case, we say that an undirected graph $G$ is a robust $(\nu,\tau)$-expander if the associated symmetric digraph is a robust $(\nu,\tau)$-out expander.
  • Proof.
    We may assume without loss of generality that $\tau\leq 1/2$ and that $\alpha\leq 1$. Let $t=\lfloor\log_{1-\tau/2}(\alpha/2)\rfloor,$ and let
    $\gamma,\nu\ll\rho_{0}\ll\rho_{1}\ll\cdots\ll\rho_{t}\ll\delta,\tau,\alpha.$
    Let $G_{0}:=G$ and $n_{0}:=n$. For each $0\leq i\leq t$, in turn, locate a vertex set $U_{i}\subseteq V(G_{i})$ satisfying $\tau n_{i}/2\leq|U_{i}|\leq(1-\tau/2)n_{i}$ and
    $\partial^{+}_{G_{i}}(U_{i})+\partial^{-}_{G_{i}}(U_{i})\leq\rho_{i}n^{2},$
    if such a set exists. In this case, define $G_{i+1}:=G_{i}[U_{i}]$ and $n_{i+1}:=|V(G_{i+1})|$. By replacing $U_{i}$ by its complement in $G_{i}$ if necessary, we can assume that
    $\frac{1}{|U_{i}|}\sum_{v\in U_{i}}w(v)\leq\frac{1}{|V(G_{i})|}\sum_{v\in V(G_{i})}w(v).$ (3)
    If such a set does not exist, then terminate this process, and define $s:=i$. If, on the other hand, we manage to find a suitable set $U_{i}$ for every $0\leq i\leq t$, then define $s:=t+1$.
    First, we observe that we cannot have $n_{s}\leq\alpha n/2$. Suppose to the contrary, and consider the minimum index $i$ for which $n_{i}\leq\alpha n/2$. Note that we also have $n_{i}\geq\tau n_{i-1}/2>\tau\alpha n/4$ by the minimality of $i$, which means
    $\partial_{G}^{+}(V(G_{i}))\leq\sum_{j=0}^{i-1}\partial^{+}_{G_{j}}(U_{j})\leq\sum_{j=0}^{i-1}\rho_{j}n^{2}\leq 2\rho_{i-1}n^{2}\leq\frac{\tau\alpha^{2}n^{2}}{8}<\frac{\alpha nn_{i}}{2}$
    since $\rho_{0}\ll\cdots\ll\rho_{i-1}\ll\tau,\alpha$. But then
    $\alpha nn_{i}\leq\sum_{v\in V(G_{i})}\deg_{G}^{+}(v)=e(V(G_{i}))+\partial^{+}_{G}(V(G_{i}))<n_{i}^{2}+\frac{\alpha nn_{i}}{2}\leq\alpha nn_{i},$
    which is clearly false. We conclude that $n_{s}>\alpha n/2$. In particular, this implies that we cannot have $s=t+1$, as $\alpha n/2<n_{s}\leq(1-\tau/2)^{s}n$, so $s<\log_{1-\tau/2}(\alpha/2)<t+1$.
    Now by our choice of $s$, we must have been unable to locate a suitable set $U_{i}$ when we reached $i=s$. Therefore, we have for every $U\subseteq V(G_{s})$ with $\tau n_{s}/2\leq|U|\leq(1-\tau/2)n_{s}$ that
    $\partial_{G_{s}}^{+}(U)+\partial^{-}_{G_{s}}(U)\geq\rho_{s}n^{2}.$
    We also have that
    $|\partial^{+}_{G_{s}}(U)-\partial^{-}_{G_{s}}(U)|$ $\leq|\partial_{G}^{+}(U)-\partial_{G}^{-}(U)|+\partial_{G}^{+}(V(G_{s}))+\partial_{G}^{-}(V(G_{s}))$
    $\leq\left|\sum_{v\in U}\deg_{G}^{+}(v)-\deg_{G}^{-}(v)\right|+\sum_{j=0}^{s-1}\rho_{j}n^{2}$
    $\leq\gamma n^{2}+2\rho_{s-1}n^{2}\leq\frac{\rho_{s}n^{2}}{2},$
    since $\gamma\ll\rho_{0}\ll\cdots\ll\rho_{s-1}\ll\rho_{s}$. Thus $\partial_{G_{s}}^{+}(U)\geq\rho_{s}n^{2}/4$, and $\partial_{G_{s}}^{-}(U)\geq\rho_{s}n^{2}/4$.
We now obtain $G^{\prime}$ from $G_{s}$ by deleting the set $Y$ of vertices $v\in V(G_{s})$ with
$\deg^{+}_{G}(v,V(G)\setminus V(G_{s}))+\deg^{-}_{G}(v,V(G)\setminus V(G_{s}))\geq\sqrt{\rho_{s-1}}n.$
We claim that $G^{\prime}$ satisfies A1, A2, and A3 with $\nu:=\rho_{s}/16$. Note that
$|Y|\sqrt{\rho_{s-1}}n\leq\partial^{+}_{G}(V(G_{s}))+\partial^{-}_{G}(V(G_{s}))\leq\sum_{j=0}^{s-1}\rho_{j}n^{2}\leq 2\rho_{s-1}n^{2},$
so $|Y|\leq 2\sqrt{\rho_{s-1}}n\leq\alpha n/4\leq n_{s}/2$ since $\rho_{s-1}\ll\alpha$. Thus every vertex $v\in V(G^{\prime})$ has in-degree at least $\deg_{G}^{-}(v)-3\sqrt{\rho_{s-1}}n\geq\deg_{G}^{-}(v)-\delta n$ and out-degree at least $\deg_{G}^{+}(v)-3\sqrt{\rho_{s-1}}n\geq\deg_{G}^{+}(v)-\delta n$ since $\rho_{s-1}\ll\delta$. This verifies A2. We also have
$|V(G^{\prime})|=n_{s}-|Y|\geq n_{s}-2\sqrt{\rho_{s-1}}n\geq n_{s}/2$
since $n_{s}>\alpha n/2$ and $\rho_{s-1}\ll\alpha$. Consequently, if $U\subseteq V(G^{\prime})$ with $\tau|V(G^{\prime})|\leq|U|\leq(1-\tau)|V(G^{\prime})|$, then $\tau n_{s}/2\leq|U|\leq(1-\tau/2)n_{s}$, so
$\partial^{+}_{G^{\prime}}(U)\geq\partial^{+}_{G_{s}}(U)-|Y|n\geq\frac{\rho_{s}n^{2}}{4}-\sqrt{\rho_{s-1}}n^{2}\geq\frac{\rho_{s}n^{2}}{8}$
since $\rho_{s-1}\ll\rho_{s}$. Then, by the definition of $RN^{+}_{\nu,G^{\prime}}(U)$ and since $\nu\ll\rho_{s}$,
$2\nu|V(G^{\prime})|^{2}\leq\frac{\rho_{s}n^{2}}{8}\leq\partial^{+}_{G^{\prime}}(U)\leq|RN^{+}_{\nu,G^{\prime}}(U)\setminus U||V(G^{\prime})|+\nu|V(G^{\prime})|^{2},$
which gives $|RN^{+}_{\nu,G^{\prime}}(U)\setminus U|\geq\nu|V(G^{\prime})|$. That is, $G^{\prime}$ is a robust $(\nu,\tau)$-out-expander, as claimed. A similar argument shows that $G^{\prime}$ is a robust $(\nu,\tau)$-in-expander as well, so A1 is satisfied. To verify A3, we note that (3) implies by induction that
$\frac{1}{n_{s}}\sum_{v\in V(G_{s})}w(v)\leq\frac{1}{n}\sum_{v\in V(G)}w(v).$
Now because $|V(G^{\prime})|\geq n_{s}/2$, we have
$\frac{1}{|V(G^{\prime})|}\sum_{v\in V(G^{\prime})}w(v)\leq\frac{2}{n_{s}}\sum_{v\in V(G_{s})}w(v)\leq\frac{2}{n}\sum_{v\in V(G)}w(v).$
This verifies A3, and the proof is complete. ∎
We proceed to prove a key technical property of dense digraphs with robust out-expansion properties. Namely, one can robustly find short rainbow paths between any two vertices whose internal vertices and colours come from fixed small random sets $V_{0}$ and $C_{0}$, respectively. This is made precise in the following lemma. Our intended application is to set $V_{0}$ and $C_{0}$ aside as a “reservoir”, and the remaining coloured directed graph $G^{\prime}$ will inherit the essential properties of $G$. Using the results of the previous section, we can find a large rainbow path forest in $G^{\prime}$, and then we can use the reservoir to connect its components into a single rainbow path in $G$ (see Theorem 6.4 and Lemma 5.1).
Lemma 4.3.
Let $n$ be a positive integer, and let $\nu,\tau,\alpha,p\leq 1$ be positive constants satisfying $\nu+\tau\leq\alpha$ and $p^{3}\nu^{2}n\geq 144\log n$. Define $\beta:=p^{3}\nu/100$.
Let $G$ be a properly edge-coloured directed graph on $n$ vertices. Suppose that $G$ is a robust $(\nu,\tau)$-out-expander, with $\delta^{\pm}(G)\geq\alpha n$. Let $V_{0},C_{0}$ be independent $p$-random subsets of $V(G),C(G)$, respectively. Then with probability at least $1-5/n$, the following all hold:
  1. B1
    $|V_{0}|\leq 2pn$;
  2. B2
    every vertex $v\in V(G)$ satisfies
    $\deg^{+}(v,V(G)\setminus V_{0};C(G)\setminus C_{0})\geq(1-3p)\deg^{+}(v);$
    $\deg^{-}(v,V(G)\setminus V_{0};C(G)\setminus C_{0})\geq(1-3p)\deg^{-}(v);$
  3. B3
    for any two distinct vertices $u,v\in V(G)$, and for any vertex subset $V_{1}\subseteq V_{0}$ and colour subset $C_{1}\subseteq C_{0}$, each of size at most $\beta n$, there exists a rainbow directed path of length at most $\nu^{-1}+1$ from $u$ to $v$ in $G$ whose internal vertices are in $V_{0}\setminus V_{1}$ and whose colours are in $C_{0}\setminus C_{1}$.
  • Proof.
    The properties B1 and B2 occur with high probability by Chernoff’s inequality (Lemma 2.1). Indeed, the probability that B1 fails is at most
    $\exp\left(-\frac{pn}{3}\right)\leq n^{-48}$
    since $pn\geq p^{3}\nu^{2}n\geq 144\log n$. As for B2, we have for every $v\in V(G)$ that each of $\deg^{+}(v)-\deg^{+}(v,V(G)\setminus V_{0};C(G)\setminus C_{0})$ and $\deg^{-}(v)-\deg^{-}(v,V_{0};C(G)\setminus C_{0})$ is a binomial random variable with expectation $(2p-p^{2})\deg^{+}(v)$ and $(2p-p^{2})\deg^{-}(v)$, respectively. Therefore, because $\delta^{\pm}(G)\geq\alpha n$, the probability that B2 fails is at most
    $2n\exp\left(-\frac{(p+p^{2})^{2}\delta^{\pm}(G)}{2(2p-p^{2})+(p+p^{2})}\right)\leq 2n\exp\left(-\frac{p\alpha n}{5}\right)\leq 2n^{-139/5}$
    since $p\alpha n\geq p^{3}\nu^{2}n\geq 144\log n$.
    In order to ensure B3, we introduce some notation. For a vertex $u\in V(G)$, define $N_{1}(u):=N_{G}^{+}(u)$, and for each $i\geq 1$, define $N_{i+1}(u):=N_{i}(u)\cup RN_{\nu,G}^{+}(N_{i}(u))$. Then $|N_{1}(u)|\geq\alpha n\geq\tau n$, so because $G$ is a robust $(\nu,\tau)$-out-expander, we have $|N_{i+1}(u)|\geq|N_{i}(u)|+\nu n$ as long as $|N_{i}(u)|\leq(1-\tau)n$. Thus there exists $i_{0}\leq\nu^{-1}$ such that $|N_{i_{0}}(u)|\geq(1-\tau)n$. As any vertex in $V(G)$ has at least $\alpha n-\tau n\geq\nu n$ in-neighbours in $N_{i_{0}}(u)$, we have that $RN_{\nu,G}^{+}(N_{i_{0}}(u))=V(G)$.
    Given $u,w\in V(G)$ with $w\in RN_{\nu/2,G}^{+}(N_{1}(u))$, let $Y_{u,w,1}$ be the number of in-neighbours $x$ of $w$ in $N_{1}(u)$ such that $x\in V_{0}$, and $C(u,x)$ and $C(x,w)$ are both in $C_{0}$. Similarly, if $w\in RN_{\nu/2,G}^{+}(N_{i}(u)\setminus N_{1}(u))$ for some $2\leq i\leq i_{0}$, let $Y_{u,w,i}$ be the number of in-neighbours $x$ of $w$ in $N_{i}(u)\setminus N_{1}(u)$ such that $x\in V_{0}$, and $C(x,w)\in C_{0}$. We will ensure that
    $Y_{u,w,1}\geq 3(\beta n+\nu^{-1})+1$ (4)
    whenever $w\in RN_{\nu/2,G}^{+}(N_{1}(u))$, and that
    $Y_{u,w,i}\geq 2(\beta n+\nu^{-1})+1$ (5)
    whenever $w\in RN_{\nu/2,G}^{+}(N_{i}(u)\setminus N_{1}(u))$.
    For $2\leq i\leq i_{0}$ and $w\in RN_{\nu/2,G}^{+}(N_{i}(u)\setminus N_{1}(u))$, $Y_{u,w,i}$ is a binomial random variable with expectation at least
    $\frac{p^{2}\nu n}{2}\geq\frac{p^{3}\nu n}{4}+\frac{p^{3}\nu n}{4}\geq 25\beta n+36\nu^{-1}\geq 2(2(\beta n+\nu^{-1})+1),$
    since $p^{3}\nu^{2}n\geq 144\log n\geq 144.$ Thus, the probability that some $Y_{u,w,i}<2(\beta n+\nu^{-1})+1$ for $i\geq 2$ is at most
    $\nu^{-1}n^{2}\exp\left(-\frac{p^{2}\nu n}{16}\right)\leq n^{-6}$
    by Chernoff’s inequality (Lemma 2.1) and a union bound. Here, we used that $\nu n\geq p^{3}\nu^{2}n\geq 144\log n\geq 1$, so $\nu^{-1}\leq n$.
    If $w\in RN_{\nu/2,G}^{+}(N_{1}(u))$, then we can greedily obtain a set $W_{u,w}\subseteq N_{G}^{-}(w)\cap N_{1}(u)$ of size at least $\nu n/6$ such that all $2|W_{u,w}|$ edges of the form $(u,x)$ or $(x,w)$ for $x\in W_{u,w}$ have distinct colours. Now $Y_{u,w,1}$ is at least the number of vertices $x\in W_{u,w}\cap V_{0}$ such that $C(u,x)$ and $C(x,w)$ are both in $C_{0}$, which is a binomial random variable with expectation at least
    $\frac{p^{3}\nu n}{6}=\frac{p^{3}\nu n}{12}+\frac{p^{3}\nu n}{12}\geq\frac{25\beta n}{3}+12\nu^{-1}\geq 2(3(\beta n+\nu^{-1})+1).$
    Thus the probability that some $Y_{u,w,1}<3(\beta n+\nu^{-1})+1$ is at most
    $n^{2}\exp\left(-\frac{p^{3}\nu n}{48}\right)\leq n^{-1}.$
    We now fix an outcome $V_{0}$ and $C_{0}$ for which B1, B2, (4), and (5) all hold (for any $u,w,i$ in the case of (4) and (5)). We will now show that whenever these conditions hold, B3 is also satisfied. Since we have shown that these conditions hold with probability at least $1-n^{-48}-2n^{-139/5}-n^{-6}-n^{-1}\geq 1-5/n$, this will complete the proof.
Fix two distinct vertices $u,v\in V(G)$, a vertex subset $V_{1}\subseteq V_{0}$, and a colour subset $C_{1}\subseteq C_{0}$, each of size at most $\beta n$. Our goal is to find a rainbow directed path $P$ from $u$ to $v$ of length at most $\nu^{-1}+1$, with $V(P)\setminus\{u,v\}\subseteq V_{0}\setminus V_{1}$ and $C(P)\subseteq C_{0}\setminus C_{1}$.
Claim.
There exist a vertex $w\in RN^{+}_{\nu/2,G}(N_{1}(u))$ and a rainbow directed path $Q$ from $w$ to $v$ of length at most $\nu^{-1}-1$, with $V(Q)\setminus\{v\}\subseteq V_{0}\setminus V_{1}$ and $C(Q)\subseteq C_{0}\setminus C_{1}$.
  • Proof.
    Recall that $v\in RN_{\nu,G}^{+}(N_{i_{0}}(u))=V(G)$. Define $v_{0}:=v$. Let $k\geq 0$ be maximal so that there exists some rainbow directed path $P_{k}:=(v_{k},v_{k-1},\ldots,v_{0})$ such that for indices $1\leq i_{k}<i_{k-1}<\cdots<i_{0}$, we have
      $v_{j}\in RN_{\nu,G}^{+}(N_{i_{j}}(u))$ for $0\leq j\leq k$;
      $V(P_{k})\setminus\{v_{0}\}\subseteq V_{0}\setminus V_{1}$;
      $C(P_{k})\subseteq C_{0}\setminus C_{1}$.
    The trivial path at $v_{0}$ satisfies all of these conditions with $k=0$, so $P_{k}$ is well-defined. We will show that the claim is satisfied with $w:=v_{k}$ and $Q:=P_{k}$. First note that $i_{1}>\cdots>i_{k}$ are $k$ distinct integers in $[1,i_{0}-1]$, so the length of $Q$ is $k\leq i_{0}-1\leq\nu^{-1}-1$. It remains to show that $w$ has at least $\nu n/2$ in-neighbours in $N_{1}(u)$.
    Suppose not. Then since $w=v_{k}\in RN^{+}_{\nu,G}(N_{i_{k}}(u))$, we must have $i_{k}\geq 2$ and $v_{k}\in RN_{\nu/2,G}^{+}(N_{i_{k}}(u)\setminus N_{1}(u))$. By (5), $v_{k}$ has at least $12(\beta n+\nu^{-1})+$ 1 in-neighbours $x\in N_{i_{k}}(u)\setminus N_{1}(u)$ such that $x\in V_{0}$ and $C(x,w)\in C_{0}$. Of these, at most $\beta n$ are in $V_{1}$, at most $\beta n$ have $C(x,w)\in C_{1}$, at most $k+1\leq\nu^{-1}$ are already in $V(P_{k})$, and at most $k\leq\nu^{-1}$ have $C(x,w)\in C(P_{k})$. This leaves at least one choice of an in-neighbour $v_{k+1}\in N_{i_{k}}(u)\setminus N_{1}(u)$ for which $P_{k+1}:=(v_{k+1},\ldots,v_{0})$ is a rainbow directed path with $V(P_{k+1})\setminus\{v_{0}\}\subseteq V_{0}\setminus V_{1}$ and $C(P_{k+1})\subseteq C_{0}\setminus C_{1}$. By definition of $N_{i_{k}}(u)\setminus N_{1}(u)$, we have that $v_{k+1}\in RN^{+}_{\nu,G}(N_{i_{k+1}}(u))$ for some $1\leq i_{k+1}<i_{k}$, contradicting the maximality of $k$.
    We conclude that $w\in RN_{\nu/2,G}^{+}(N_{1}(u))$. This proves the claim. ∎
Let $w$ and $Q$ be as in the claim. If $u\in V(Q)$, then we can take $P$ to be the subpath of $Q$ from $u$ to $v$. If not, then by (4), $w$ has at least $13(\beta n+\nu^{-1})+$ 1 in-neighbours $x\in N_{1}(u)$ such that $x\in V_{0}$, and $C(u,x),C(x,w)\in C_{0}$. Of these, at most $\beta n$ are in $V_{1}$, at most $2\beta n$ have $C(u,x)\in C_{1}$ or $C(x,w)\in C_{1}$, at most $|V(Q)|\leq\nu^{-1}$ are already in $V(Q)$, and at most $2|C(Q)|\leq 2\nu^{-1}$ have $C(u,x)\in C(Q)$ or $C(x,w)\in C(Q)$. This leaves at least one choice for $x$ in which $P:=(u,x,Q)$ is a rainbow directed path from $u$ to $v$ with $V(P)\setminus\{u,v\}\subseteq V_{0}\setminus V_{1}$ and $C(P)\subseteq C_{0}\setminus C_{1}$. Because the length of $P$ is at most $\nu^{-1}+1$, the proof is complete. ∎
5 Long rainbow paths in dense digraphs
In this section we prove Theorem 1.6 and Theorem 1.4(c). The rough idea of both proofs is as follows. Dense regular digraphs (and more generally Eulerian digraphs) without good expansion properties have sparse cuts, so we can always zoom into an expanding subgraph with essentially the same degrees as the original digraph (as in Lemma 4.2). On the other hand, dense digraphs with good expansion properties (including all sufficiently dense Cayley graphs on $\mathbb{Z}_{p}$ with $p$ prime, as we will see) are well-connected, so we can build a rainbow path by first finding a large rainbow path forest with few components via Lemma 3.2, and then connecting the components together via short rainbow paths via Lemma 4.3.
We begin with the case where our dense digraph already has robust expansion properties.
Lemma 5.1.
For any $\varepsilon>0$, there exists a positive integer $n_{0}$ such that the following holds for all $n\geq n_{0}$. Let $\nu,\tau,\alpha>0$ satisfying $\nu+\tau\leq\alpha\leq 1$ and $\nu\geq n^{-1/2}\log n$. Let $G$ be a properly edge-coloured directed graph on $n$ vertices. Suppose that $G$ is a robust $(\nu,\tau)$-out-expander with $\delta^{\pm}(G)\geq\alpha n$. Then $G$ contains a rainbow directed path of length at least $\left(1-\varepsilon\right)\delta^{\pm}(G)$.
  • Proof.
    We may assume $\varepsilon\leq 1$. Let $d:=\delta^{\pm}(G)$. Set $p:=\varepsilon/6$ and $\beta:=p^{3}\nu/100$. Since $\nu\geq n^{-1/2}\log n$ and $1/n\ll\varepsilon$, we have
    $p^{3}\nu^{2}n\geq\frac{\varepsilon^{3}\log^{2}n}{216}\geq 144\log n,$
    so by Lemma 4.3, there exist sets $V_{0}\subseteq V(G)$ and $C_{0}\subseteq C(G)$ satisfying B1B3. Let $G^{\prime}:=G[V(G)\setminus V_{0};C(G)\setminus C_{0}]$. By B2, the minimum in-degree of $G^{\prime}$ is at least
    $(1-3p)d\geq(1-\varepsilon/2)\alpha n\geq\frac{\nu n}{2}\geq\frac{\sqrt{n}\log n}{2}\geq 9(\varepsilon/2)^{-3}$
    since $\alpha\geq\nu\geq n^{-1/2}\log n$, and $1/n\ll\varepsilon\leq 1$. Therefore, by Lemma 3.2 applied to $G^{\prime}$ with $\varepsilon/2$ in place of $\varepsilon$, $G^{\prime}$ has a rainbow path forest $\mathcal{P}$ with at most $36\varepsilon^{-2}$ components and at least $(1-\varepsilon/2)(1-3p)d\geq(1-\varepsilon)d$ edges.
    Let $P_{1},\ldots,P_{m}$ be the components of $\mathcal{P}$, where $m\leq 36\varepsilon^{-2}$. For each $1\leq i\leq m$, let $u_{i}$ denote the vertex of $P_{i}$ with in-degree 00, and let $w_{i}$ denote the vertex of $P_{i}$ with out-degree 00. Now for each $11\leq i\leq m-$ 1, in turn, find a path $Q_{i}$ from $w_{i}$ to $u_{i+1}$ of length at most $\nu^{-1}+1$ whose internal vertices are in $V_{0}\setminus\bigcup_{j<i}V(Q_{j})$ and whose colours are in $C_{0}\setminus\bigcup_{j<i}C(Q_{j})$. Such paths exists by B3, as for each $i$, the sizes of $\bigcup_{j<i}V(Q_{j})$ and $\bigcup_{j<i}C(Q_{j})$ never exceed
    $m\left(\nu^{-1}+2\right)\leq 36\varepsilon^{-2}\left(\frac{\sqrt{n}}{\log n}+2\right)\leq\frac{\varepsilon^{3}\sqrt{n}\log n}{21600}\leq\frac{p^{3}\nu n}{100}=\beta n$
    since $\nu\geq n^{-1/2}\log n$ and $1/n\ll\varepsilon$. Now $\bigcup_{i=1}^{m}P_{i}\cup\bigcup_{i=1}^{m-1}Q_{i}$ is a rainbow path in $G$ of length at least $(1-\varepsilon)d$, as required. ∎
5.1 Dense digraphs
We now prove Theorem 1.6 in the following stronger form. Here we relax the condition that the digraph $G$ is $d$-regular with $d=\Omega(n)$ to the condition that $G$ is almost-Eulerian (i.e., each vertex has in-degree within $o(n)$ of its out-degree) with $\delta^{\pm}(G)=\Omega(n)$.
Theorem 5.2.
For any $\varepsilon,\alpha>0$, there exist $n_{0}\in\mathbb{N}$ and $\gamma>0$ such that the following holds for all $n\geq n_{0}$. Let $G$ be a properly edge-coloured directed graph on $n$ vertices with $d:=\delta^{\pm}(G)\geq\alpha n$. Suppose that every vertex $v\in V(G)$ satisfies $|\deg^{+}(v)-\deg^{-}(v)|\leq\gamma n$. Then $G$ has a rainbow directed path of length at least $(1-\varepsilon)d$.
  • Proof.
    We may assume without loss of generality that $\varepsilon\leq 1$. Let $\delta:=\varepsilon\alpha/2$, and let $\tau,\nu,\gamma>0$ be such that $1/n\ll\gamma,\nu\ll\tau\ll\varepsilon,\alpha$. Let $G^{\prime}$ be as in Lemma 4.2, satisfying A1 and A2 (we don’t need A3 for this application). By A2, we have $\delta^{\pm}(G^{\prime})\geq d-\varepsilon\alpha n/2\geq(1-\varepsilon/2)d\geq\alpha n/2$. We also have $\nu\geq|V(G^{\prime})|^{-1/2}\log|V(G^{\prime})|$ and $\nu+\tau\leq(1-\varepsilon/2)\alpha$ since $|V(G^{\prime})|\geq\delta^{\pm}(G^{\prime})\geq\alpha n/2$ and $1/n\ll\nu,\tau\ll\alpha,\varepsilon\leq 1$. Now by Lemma 5.1 with $\varepsilon/2$, $|V(G^{\prime})|$, $(1-\varepsilon/2)\alpha$, $G^{\prime}$, in place of $\varepsilon$, $n$, $\alpha$, $G$, respectively, $G^{\prime}$ contains a rainbow directed path of length at least
    $\left(1-\varepsilon/2\right)\delta^{\pm}(G^{\prime})\geq(1-\varepsilon)d.$
    This completes the proof.∎
Remark. Essentially the same proof also guarantees the existence of a directed cycle of length $d-o(d)$ in both Lemma 5.1 and Theorem 5.2. Moreover, our proof of Theorem 5.2 can be optimized so that the conclusion holds whenever $\delta^{\pm}(G)\geq 100n\log\log n/\log n$, but to keep our computations neat, we prove this slightly weaker version assuming $\delta^{\pm}(G)=\Omega(n)$.
5.2 Graham’s original problem for $\mathbb{Z}_{p}$
In this section we prove Theorem 1.4(c) as an immediate consequence of Lemma 5.1. We will use a classical theorem of Pollard from 1970 [57] generalising the classical Cauchy-Davenport inequality to show that any dense Cayley graph on $\mathbb{Z}_{n}$ for prime $n$ is a robust out-expander [^3] Throughout this section, we denote the (prime) order of the group by $n$ instead of $p$ in order to keep consistency with the rest of the paper.. To state it, we first require some notation. Given two subsets $A,B$ of an abelian group $\Gamma$ and an element $x\in\Gamma$, we denote by $r_{A,B}(x)$ the number of $(a,b)\in A\times B$ such that $x=a+b$. Let us define the $t$-representable sum of $A$ and $B$ as
$A+_{t}B:=\{x\in\Gamma\mid r_{A,B}(x)\geq t\}.$
We are now ready to state Pollard’s theorem.
Theorem 5.3.
Let $n$ be a prime and $A,B$ be nonempty subsets of $\mathbb{Z}_{n}$. Then for any $t\leq\min\{|A|,|B|\}$ we have
$\sum_{i=1}^{t}|A+_{i}B|\geq t\cdot\min\{n,|A|+|B|-t\}.$
We now prove Theorem 1.4(c). We restate the theorem in terms of rainbow paths in Cayley graphs.
Theorem 5.4.
For any $\varepsilon>0$, there exists $n_{0}$ such that the following holds for all primes $n\geq n_{0}$. Let $S\subseteq\mathbb{Z}_{n}\setminus\{0\}$ of size $|S|\geq 4n^{3/4}\sqrt{\log n}$. Then $\mathrm{Cay}(\mathbb{Z}_{n},S)$ contains a rainbow path of length $(1-\varepsilon)|S|$.
  • Proof.
    Let $\varepsilon>0$, and assume that $n$ is a sufficiently large prime. Let $S\subseteq\mathbb{Z}_{n}\setminus\{0\}$ of size $d:=|S|\geq 4n^{3/4}\sqrt{\log n}$, and let $G:=\mathrm{Cay}(\mathbb{Z}_{n},S)$. Also define $\nu:=d^{2}/(8n^{2})$, and let $\tau:=d/(2n)$. We will first show that $G$ is a robust $(\nu,\tau)$-out-expander. Let $U\subseteq V(G)$ with $\tau n\leq|U|\leq(1-\tau)n$, and let $t:=\lceil d/2\rceil$. Note that $|S|,|U|\geq t$, so we can apply Theorem 5.3 with $U$, $S$ in place of $A$, $B$ to conclude that
    $t(|U|+|S|-t)=t\cdot\min\left\{n,|U|+|S|-t\right\}\leq\sum_{i=1}^{t}|U+_{i}S|=\sum_{x\in\mathbb{Z}_{n}}\min\left\{t,r_{U,S}(x)\right\}=\sum_{x\in V(G)}\min\left\{t,\deg^{-}_{G}(x,U)\right\},$
    where the first equality used that $|U|\leq(1-\tau)n$, so $|U|+|S|-t\leq|U|+d/2\leq n$. Defining $N:=RN^{+}_{\nu,G}(U)\setminus U$, we can estimate the rightmost sum by
    $\sum_{x\in V(G)}\min\left\{t,\deg^{-}_{G}(x,U)\right\}\leq t|U|+t\left|N\right|+\nu n(n-|U|-|N|)\leq t|U|+t|N|+\nu n^{2},$
    which gives
    $|N|\geq\frac{t(|S|-t)-\nu n^{2}}{t}\geq\left\lfloor d/2\right\rfloor-d/4\geq d^{2}/(8n)=\nu n$
    since $4\leq d\leq n$. Hence $G$ is a robust $(\nu,\tau)$-out-expander.
    Now set $\alpha:=d/n$ so $\delta^{\pm}(G)=d=\alpha n$, and note that $\nu+\tau\leq\alpha$ and $\nu=d^{2}/(8n^{2})>n^{-1/2}\log n$. By Lemma 5.1, $G$ contains a rainbow directed path of length at least $(1-\varepsilon)d$. ∎
We note that there has been plenty of work on generalising Pollard’s Theorem to other abelian groups. One can use a generalization due to Green and Rusza [31] (see also [32]) to extend Theorem 5.4 to any abelian group $\Gamma$ but with the density requirement depending on the size of the largest proper subgroup of $\Gamma$.
6 Long rainbow paths in regular graphs
In this section, we prove Theorem 1.5, which says that we can find a long rainbow path in any regular properly edge-coloured (undirected) graph. The proof is split into two parts. In the first part (contained in Lemma 6.3), we show that if the graph expands in a particular way, then we can iteratively grow our desired long rainbow path from a single vertex. In the second part, we show that when this expansion property fails, then we can reduce the problem to a dense subgraph and apply the tools of Sections 3 and 4.
We start with a definition of an auxiliary object called a mop which will prove very useful in the proof.
Definition 6.1.
Let $G$ be a properly edge-coloured graph. For non-negative integers $\ell,s,t$, an $(\ell,s,t)$-mop in $G$ is a quadruple $(P,v,U,\{Q_{u}\}_{u\in U})$ such that $P\subseteq G$ is a rainbow path of length $\ell$ with $v$ as an endpoint, $U\subseteq V(G)$ is a set of $t$ vertices, and for each $u\in U$, $Q_{u}\subseteq G$ is a path from $v$ to $u$ of length at most $s$ such that $P\cup Q_{u}$ is a rainbow path. We call $P$ the handle of the mop, the vertices in $U$ the ends, and the $Q_{u}$’s the strands.
To give some intuition behind this definition, recall the standard greedy argument for finding long paths in graphs. We first take the longest path and argue that if it was too short, then the end-vertex could not send all of its neighbours to the path and hence one can use one neighbour to extend the path, contradicting maximality. The issue with this argument in the rainbow setting is that in addition to not sending edges to the already built path we must also not reuse colours appearing on the path. The way around this is that instead of looking at a single longest path we look for an object which consists of a long rainbow path (the handle) and many short rainbow continuations (strands) through which we can reach many vertices (ends). The major benefit of this approach is that now we have many potential end-vertices of a long rainbow path instead of just one. This allows one to easily extend one of these paths by a single vertex, but in order to iterate the argument, this is not good enough; we want to preserve the mop structure as well. The following definition captures a weak expansion property we require from the ends in order to be able to perform the iteration.
Definition 6.2.
We say that an $(\ell,s,t)$-mop $(P,v,U,\{Q_{u}\}_{u\in U})$ in a properly edge-coloured graph $G$ is $(d,\gamma)$-leaky if $e(U,V(G)\setminus U;C(G)\setminus C(P))\geq\gamma(d-\ell)|U|$. We say that $G$ is a $(d,s_{0},\gamma,\alpha)$-leaky mop expander if every $(\ell,s,t)$-mop in $G$ with $s\leq s_{0}$ and $t\leq\alpha^{-1}(d-\ell)$ is $(d,\gamma)$-leaky.
We are now ready to prove that if $G$ has this weak expansion property, then we can build our desired long path. The basic idea is that a leaky $(\ell,s,t)$-mop with the number $t$ of ends large enough compared to the length $\ell$ of the handle can be made significantly larger (increasing $t$) by increasing the maximum length $s$ of the strands by one. If $t$ is very large compared to $\ell$, then we can extend the handle to build a longer mop (increasing $\ell$) by incorporating a strand whose end is sufficiently “far” from the handle in an appropriate sense. In a $d$-regular leaky mop expander, this can be iterated until the handle has length $d-o(d)$.
Lemma 6.3.
Let $1/d\ll\alpha\ll 1/s_{0}\ll\gamma,\varepsilon,1/K\leq 1$, and let $G$ be a properly edge-coloured graph with $d=\delta(G)\leq\Delta(G)\leq Kd$. If $G$ is a $(d,s_{0},\gamma,\alpha)$-leaky mop expander, then $G$ contains a rainbow path of length at least $(1-\varepsilon)d$.
  • Proof.
    Introduce a further parameter $s_{1}$ so that $1/s_{0}\ll 1/s_{1}\ll\gamma,\varepsilon,1/K$. We consider $(\ell,s,t)$-mops in $G$ satisfying the following properties:
    1. (a)
      $1\leq s\leq s_{0}$, and if $\ell>0$, then $s\geq s_{1}$;
    2. (b)
      ${t\geq\left(1+\frac{\gamma\varepsilon}{2K}\right)^{s-1}\frac{\varepsilon d}{2}}.$
    Clearly, taking any vertex $v$ and $d$ of its neighbours gives a $(0,1,d)$-mop, so such mops exist. Now consider such a mop $(P,v,U,\{Q_{u}\}_{u\in U})$ which maximises $\ell$, then $s$. Define $C^{\prime}:=C(G)\setminus C(P)$ and $d^{\prime}:=d-\ell$. We will show that $\ell\geq(1-\varepsilon)d$, and so $P$ is our desired rainbow path. Suppose that this is not the case, so that $d^{\prime}\geq\varepsilon d$. We consider two cases and derive a contradiction in each one.
    If $s<s_{0}$, then $(P,v,U,\{Q_{u}\}_{u\in U})$ is also an $(\ell,s+1,t)$-mop satisfying (a), so by the maximality of $s$, (b) must fail, and we must have $t\leq\left(1+\frac{\gamma\varepsilon}{2K}\right)^{s}\frac{\varepsilon d}{2}\leq\left(\frac{3}{2}\right)^{s}\frac{\varepsilon d}{2}\leq\alpha^{-1}d^{\prime}$ since $\gamma,\varepsilon,1/K\leq 1$ and $\alpha\ll 1/s_{0},\varepsilon$. Therefore, since $G$ is a $(d,s_{0},\gamma,\alpha)$-leaky mop expander, we have
    $e(U,V(G)\setminus U;C^{\prime})\geq\gamma d^{\prime}|U|.$ (6)
    We will count these edges directly to show that this cannot happen. Let $U^{\prime}$ be the set of vertices $u^{\prime}\in V(G)\setminus U$ with some neighbour $u\in U$ such that $u^{\prime}\notin V(P\cup Q_{u})$ and $C(uu^{\prime})\in C^{\prime}\setminus C(Q_{u})$, (so that for $\tilde{Q}_{u^{\prime}}:=Q_{u}\cup\{uu^{\prime}\}$, we have that $P\cup\tilde{Q}_{u^{\prime}}$ is a rainbow path). Then $(P,v,U\cup U^{\prime},\{Q_{u}\}_{u\in U}\cup\{\tilde{Q}_{u^{\prime}}\}_{u^{\prime}\in U^{\prime}}\})$ is an $(\ell,s+1,|U\cup U^{\prime}|)$-mop, so again, by the maximality of $s$, we must have $|U\cup U^{\prime}|<\left(1+\frac{\gamma\varepsilon}{2K}\right)^{s}\frac{\varepsilon d}{2}$, which implies $|U^{\prime}|<\frac{\gamma\varepsilon}{2K}|U|$. Also, the $C^{\prime}$-coloured edges $uu^{\prime}$ with $u\in U$ and $u^{\prime}\in V(G)\setminus U$ all have $u^{\prime}\in U^{\prime}\cup V(P\cup Q_{u})$ or $C(uu^{\prime})\in C(Q_{u})$. Of these, at most $\Delta(G)|U^{\prime}|$ have $u^{\prime}\in U^{\prime}$, at most $\Delta(G)\ell$ have $u^{\prime}\in V(P)\setminus\{v\}$, at most $s|U|$ have $u^{\prime}\in V(Q_{u})$, and at most $s|U|$ have $C(uu^{\prime})\in C(Q_{u})$. This gives
    $e(U,V(G)\setminus U;C^{\prime})\leq\Delta(G)|U^{\prime}|+\Delta(G)\ell+2s|U|.$
    Since $d^{\prime}\geq\varepsilon d$ and $\Delta(G)\leq Kd$, we have $\Delta(G)|U^{\prime}|<\gamma d^{\prime}|U|/2$. Since $1/d\ll 1/s_{0},\gamma,\varepsilon$, we have $42s|U|\leq\gamma d^{\prime}|U|/$ 4. If $\ell>0$, then $s\geq s_{1}$, and so $|U|=t\geq(1+\gamma\varepsilon/(2K))^{s_{1}-1}\varepsilon d/2\geq 4\gamma^{-1}\varepsilon^{-1}Kd$ since $1/s_{1}\ll\gamma,\varepsilon,1/K$. Therefore, $Kd\ell\leq Kd^{2}\leq\gamma\varepsilon d|U|/4\leq\gamma d^{\prime}|U|/4$. If $\ell=0$, then $Kd\ell=0\leq\gamma d^{\prime}|U|/4$ as well. Hence $e(U,V(G)\setminus U;C^{\prime})<\gamma d^{\prime}|U|$, contradicting (6).
    Now suppose that $s=s_{0}$. In this case, we will attempt to construct an $(\ell^{\prime},s_{1},t^{\prime})$-mop satisfying $\ell^{\prime}>\ell$ and $t^{\prime}\geq(1+\gamma\varepsilon/(2K))^{s_{1}-1}\varepsilon d/2$ as follows. Define $B_{0}:=V(P)$, and for each $1\leq i\leq s_{1}$, define
    $B_{i}:=B_{i-1}\cup\{x\in V(G):\deg(x,B_{i-1};C^{\prime})\geq\gamma d^{\prime}/4\}.$
    Now because $\Delta(G)\leq\varepsilon^{-1}Kd^{\prime}$, we have $|B_{i}|\leq|B_{i-1}|+|B_{i-1}|\Delta(G)/(\gamma d^{\prime}/4)\leq(1+4\gamma^{-1}\varepsilon^{-1}K)|B_{i-1}|$. In particular,
    $|B_{s_{1}}|\leq\left(1+4\gamma^{-1}\varepsilon^{-1}K\right)^{s_{1}}d<\left(1+\frac{\gamma\varepsilon}{2K}\right)^{s_{0}-1}\varepsilon d/2\leq t=|U|$
    since $1/s_{0}\ll 1/s_{1},\gamma,\varepsilon,1/K$, so there exists some vertex $u\in U\setminus B_{s_{1}}$. For each $1\leq i\leq s_{1}$, let $W_{i}$ be the set of vertices $w\in V(G)\setminus B_{s_{1}-i}$ reachable from $u$ by a rainbow path $\tilde{Q}_{w,i}$ from $u$ to $w$ of length at most $i$ such that $P\cup Q_{u}\cup\tilde{Q}_{w,i}$ is also a rainbow path. By construction, we have for each $i$ that $(P\cup Q_{u},u,W_{i},\{\tilde{Q}_{w,i}\}_{w\in W_{i}})$ is a $(|E(P\cup Q_{u})|,i,|W_{i}|)$-mop. Note that $u\neq v$ since $v\in V(P)\subseteq B_{s_{1}}$, so $|E(P\cup Q_{u})|>\ell$. By the maximality of $\ell$, we must have $|W_{s_{1}}|<\left(1+\frac{\gamma\varepsilon}{2K}\right)^{s_{1}-1}\frac{\varepsilon d}{2}$. Let $1\leq i\leq s_{1}$ be the minimum index for which $|W_{i}|<\left(1+\frac{\gamma\varepsilon}{2K}\right)^{i-1}\varepsilon d/2$.
    Note that $|W_{1}|\geq\varepsilon d/2$, so $i\geq 2$. Indeed, $u$ has at least $d^{\prime}$ neighbours $w$ with $C(uw)\in C^{\prime}$. At most $\gamma d^{\prime}/4$ of these are in $B_{s_{1}-1}$ since $u\notin B_{s_{1}}$. At most $s_{0}$ are in $V(Q_{u})$, and at most $s_{0}$ have $C(uw)\in C(Q_{u})$. This gives $|W_{1}|\geq(1-\gamma/4)d^{\prime}-2s_{0}\geq 3d^{\prime}/4-2s_{0}\geq\varepsilon d/2$ since $\gamma\leq 1$, $d^{\prime}\geq\varepsilon d$, and $1/d\ll 1/s_{0},\varepsilon$.
    Consider the mop $(P\cup Q_{u},u,W_{i-1},\{\tilde{Q}_{w,i-1}\}_{w\in W_{i-1}})$. Define $C^{\prime\prime}:=C(G)\setminus C(P\cup Q_{u})$ and $d^{\prime\prime}:=d-|E(P\cup Q_{u})|$. Note that
    $|W_{i-1}|\leq|W_{i}|<\left(1+\frac{\gamma\varepsilon}{2K}\right)^{i-1}\frac{\varepsilon d}{2}\leq\alpha^{-1}d^{\prime\prime},$
    so because $G$ is a $(d,s_{0},\gamma,\alpha)$-leaky mop expander, we have that
    $e(W_{i-1},V(G)\setminus W_{i-1};C^{\prime\prime})\geq\gamma d^{\prime\prime}|W_{i-1}|.$ (7)
    As before, we count $C^{\prime\prime}$-coloured edges from $W_{i-1}$ to $V(G)\setminus W_{i-1}$ directly to show that this cannot happen. Any $C^{\prime\prime}$-coloured edge $wx$ with $w\in W_{i-1}$ and $x\in V(G)\setminus W_{i-1}$ must have $x\in(W_{i}\setminus W_{i-1})\cup B_{s_{1}-i}\cup V(Q_{u}\cup\tilde{Q}_{w,i-1})$ or $C(wx)\in C(\tilde{Q}_{w,i-1})$. By the minimality of $i$, we have
    $|W_{i}\setminus W_{i-1}|<\frac{\gamma\varepsilon|W_{i-1}|}{2K},$
    and so
    $e(W_{i-1},W_{i}\setminus W_{i-1};C^{\prime\prime})\leq\Delta(G)|W_{i}\setminus W_{i-1}|\leq\frac{\gamma d^{\prime}|W_{i-1}|}{2}.$
    Also, since $W_{i-1}\cap B_{s_{1}-i+1}=\emptyset$, we have
    $e(W_{i-1},B_{s_{1}-i};C^{\prime\prime})\leq\frac{\gamma d^{\prime}|W_{i-1}|}{4}.$
    The number of $C^{\prime\prime}$-coloured edges $wx$ with $w\in W_{i-1}$ and $x\in V(Q_{u}\cup\tilde{Q}_{w,i-1})$ or $C(wx)\in C(\tilde{Q}_{w,i-1})$ is at most $(s_{0}+2s_{1})|W_{i-1}|\leq\gamma d^{\prime}|W_{i-1}|/8$ since $d^{\prime}\geq\varepsilon d$ and $1/d\ll 1/s_{0},1/s_{1},\gamma,\varepsilon$. Altogether, we have
    $e(W_{i-1},V(G)\setminus W_{i-1};C^{\prime\prime})\leq\frac{7\gamma d^{\prime}|W_{i-1}|}{8}\leq\frac{7\gamma d^{\prime\prime}|W_{i-1}|}{8}+\frac{7\gamma s_{0}|W_{i-1}|}{8}<\gamma d^{\prime\prime}|W_{i-1}|,$
    since $1/d\ll 1/s_{0},\varepsilon$, contradicting (7).
    We conclude that $\ell\geq(1-\varepsilon)d$, and the proof is complete. ∎
We are now ready to prove Theorem 1.5. In fact, we prove the following slightly more general result for almost-regular graphs. The proof begins by identifying a mop which witnesses that the hypotheses of Lemma 6.3 fail, as otherwise we would be done. As the set of ends of the mop does not send many edges outside, we may find a robustly expanding subgraph inside it. We may then use the results of Sections 3 and 4 to find a rainbow path within this robust expander which combines with the handle of the mop to give the desired rainbow path.
Theorem 6.4.
For any $\varepsilon,K>0$, there exists $d_{0}$ such that the following holds for all $d\geq d_{0}$. Let $G$ be a properly edge-coloured graph with $d=\delta(G)\leq\Delta(G)\leq Kd$. Then $G$ has a rainbow path of length at least $(1-\varepsilon)d$.
  • Proof.
    We may assume that $\varepsilon,1/K\ll 1$. Let $1/d\ll\nu\ll\tau\ll\alpha\ll 1/s_{0}\ll\gamma\ll\varepsilon,1/K\ll 1$.
    Suppose towards a contradiction that $G$ contains no rainbow path of length at least $(1-\varepsilon)d$.
    Claim.
    There exists an $(\ell,s,n^{\prime})$-mop $(P,v,U,\{Q_{u}\}_{u\in U})$ satisfying the following properties with $d^{\prime}:=d-\ell$ and $C^{\prime}:=C(G)\setminus C(P)$:
    1. C1
      $\ell\leq(1-\varepsilon)d$, $s\leq s_{0}$, and $n^{\prime}=|U|\leq\alpha^{-1}d^{\prime}$;
    2. C2
      $e(U,U^{c};C^{\prime})<3\gamma d^{\prime}n^{\prime}$;
    3. C3
      the graph $G^{\prime}:=G[U;C^{\prime}]$ is a $(\nu,\tau)$-robust expander;
    4. C4
      $\delta(G^{\prime})\geq(1-3\gamma)d^{\prime}/2$.
  • Proof.
    By Lemma 6.3, $G$ is not a $(d,s_{0},\gamma,\alpha)$-leaky mop expander. Fix an $(\ell,s,t)$-mop $(P,v,U_{0},\{Q_{u}\}_{u\in U_{0}})$ with $s\leq s_{0}$, $t=|U_{0}|\leq\alpha^{-1}d^{\prime}$, and $e(U_{0},U_{0}^{c};C^{\prime})<\gamma d^{\prime}|U_{0}|$, where $C^{\prime}:=C(G)\setminus C(P)$ and $d^{\prime}:=d-\ell$. We have assumed that $G$ has no rainbow path of length at least $(1-\varepsilon)d$, so we have $\ell\leq(1-\varepsilon)d$ as well. Choose $U_{1}\subseteq U_{0}$ minimal with $e(U_{1},U_{1}^{c};C^{\prime})<\gamma d^{\prime}|U_{1}|$. This guarantees that
    $\delta(G[U_{1};C^{\prime}])\geq(1-\gamma)d^{\prime}/2\geq\alpha t/4\geq\alpha|U_{1}|/4.$
    Indeed, if there existed a vertex $u\in U_{1}$ whose degree in the graph $G[U_{1};C^{\prime}]$ was less than $(1-\gamma)d^{\prime}/2$, then $W:=U_{1}\setminus\{u\}$ would satisfy
    $e(W,W^{c};C^{\prime})$ $=e(U_{1},U_{1}^{c};C^{\prime})-\deg(u,U_{1}^{c};C^{\prime})+\deg(u,U_{1};C^{\prime})$
    $=e(U_{1},U_{1}^{c};C^{\prime})-\deg(u;C^{\prime})+2\deg(u,U_{1};C^{\prime})$
    $<\gamma d^{\prime}|U_{1}|-d^{\prime}+(1-\gamma)d^{\prime}$
    $=\gamma d^{\prime}|W|,$
    which would be a contradiction to the minimality of $U_{1}$. Now, recalling that $\nu\ll\tau,\alpha,\gamma$, we can apply Lemma 4.2 to the symmetric digraph associated to $G[U_{1};C^{\prime}]$ with $\gamma\alpha$, $\alpha/4$ in place of $\delta,\alpha$ and with $w(v):=\deg(v,U_{1}^{c};C^{\prime})$ to obtain a nonempty induced subgraph $G^{\prime}\subseteq G[U_{1};C^{\prime}]$ which satisfies A1-A3; that is, $G^{\prime}$ is a robust $(\nu,\tau)$-expander, every vertex $v\in U:=V(G^{\prime})$ satisfies
    $\deg_{G^{\prime}}(v)\geq\deg_{G[U_{1};C^{\prime}]}(v)-\gamma\alpha|U_{1}|\geq(1-\gamma)d^{\prime}/2-\gamma d^{\prime}=(1-3\gamma)d^{\prime}/2,$ (8)
    and
    $e(U,U_{1}^{c};C^{\prime})=\sum_{v\in U}w(v)\leq 2\frac{|U|}{|U_{1}|}\sum_{v\in U_{1}}w(v)=2\frac{|U|}{|U_{1}|}e(U_{1},U_{1}^{c};C^{\prime})<2\gamma d^{\prime}|U|.$
    Now, defining $n^{\prime}:=|U|\leq t$, the $(\ell,s,n^{\prime})$-mop $(P,v,U,\{Q_{u}\}_{u\in U})$ satisfies C1, C3, and C4. We now verify C2. Indeed, any vertex $u\in U$ sends at most $\deg_{G[U_{1};C^{\prime}]}(u)-\deg_{G^{\prime}}(u)\leq\gamma\alpha|U_{1}|\leq\gamma d^{\prime}$ edges to $U_{1}\setminus U$ in $C^{\prime}$ colours by (8), so
    $e(U,U^{c};C^{\prime})=e(U,U_{1}^{c};C^{\prime})+e(U,U_{1}\setminus U;C^{\prime})<2\gamma d^{\prime}|U|+\gamma d^{\prime}|U|=3\gamma d^{\prime}n^{\prime}.$
    Thus C2 holds as well. This proves the claim. ∎
Let $p:=\gamma$ and $\beta:=p^{3}\nu/100$. Since $n^{\prime}\geq(1-3\gamma)d^{\prime}/2\geq\varepsilon d/4$ by C1 and C4, and $1/d\ll\nu,\gamma,\varepsilon$, we have $p^{3}\nu^{2}n^{\prime}\geq 144\log n^{\prime}$. We also have $\delta(G^{\prime})\geq(1-3\gamma)d^{\prime}/2\geq\alpha n^{\prime}/4$ by C1 and C4, and $\nu+\tau\leq\alpha/4$ since $\nu\leq\tau\ll\alpha$, so we can apply Lemma 4.3 to the symmetric digraph associated to $G^{\prime}$ with $\alpha/4$, $n^{\prime}$ in place of $\alpha$, $n$, respectively, to find $V_{0}\subseteq V(G^{\prime})$ and $C_{0}\subseteq C^{\prime}$ satisfying B1-B3. Set $G^{\prime\prime}:=G^{\prime}[V(G^{\prime})\setminus V_{0};C^{\prime}\setminus C_{0}]$. Note that by B1 and B2,
$n^{\prime\prime}:=|V(G^{\prime\prime})|\geq(1-2\gamma)n^{\prime}\quad\quad\text{ and }\quad\quad\deg_{G^{\prime\prime}}(v)\geq(1-3\gamma)\deg_{G^{\prime}}(v)\>\>\text{for every }v\in V(G^{\prime\prime}).$ (9)
Claim.
There exists a rainbow path forest $\mathcal{P}$ in $G^{\prime\prime}$ with at most $144\varepsilon^{-2}$ components and at least $(1-\varepsilon/2)d^{\prime}$ edges.
  • Proof.
    We consider two cases. If $Kd\leq\varepsilon n^{\prime\prime}/16$, then $G^{\prime\prime}$ has maximum degree at most $\varepsilon n^{\prime\prime}/16$ and average degree at least
    $\frac{1}{n^{\prime\prime}}\sum_{v\in V(G^{\prime\prime})}\deg_{G^{\prime\prime}}(v)$ $\geq\frac{1}{n^{\prime\prime}}\sum_{v\in V(G^{\prime\prime})}(1-3\gamma)\deg_{G^{\prime}}(v)$
    $=\frac{1}{n^{\prime\prime}}\sum_{v\in V(G^{\prime\prime})}(1-3\gamma)(\deg_{G}(v;C^{\prime})-\deg_{G}(v,U^{c};C^{\prime}))$
    $\geq(1-3\gamma)d^{\prime}-\frac{(1-3\gamma)(3\gamma d^{\prime}n^{\prime})}{n^{\prime\prime}}$
    $\geq(1-3\gamma)d^{\prime}-\frac{(1-3\gamma)(3\gamma d^{\prime})}{1-2\gamma}\geq(1-6\gamma)d^{\prime}$
    by (9), C2, and the fact that every vertex $v\in V(G)$ satisfies $\deg_{G}(v,C^{\prime})\geq d-|C(P)|=d^{\prime}$. Thus we can apply Lemma 3.1 to $G^{\prime\prime}$ with $\varepsilon/4$, $(1-6\gamma)d^{\prime}$ in place of $\varepsilon$, $d$ to obtain a rainbow path forest with at most $\lceil\log_{8/7}(\varepsilon/4)^{-1}\rceil\leq 144\varepsilon^{-2}$ components and at least $(1-\varepsilon/4)(1-6\gamma)d^{\prime}\geq(1-\varepsilon/2)d^{\prime}$ edges, as desired.
    On the other hand, if $Kd\geq\varepsilon n^{\prime\prime}/16$, then we can find a subgraph $G^{\prime\prime\prime}$ of $G^{\prime\prime}$ with minimum degree close to $d^{\prime}$ as follows. Let $S$ be the set of vertices $u\in V(G^{\prime\prime})$ with $\deg_{G}(u,U^{c};C^{\prime})\geq\sqrt{3\gamma d^{\prime}n^{\prime}}$. Then, because $V(G^{\prime\prime})\subseteq U$ and $e_{G}(U,U^{c};C^{\prime})\leq 3\gamma d^{\prime}n^{\prime}$ by C2, we must have that $|S|\leq\sqrt{3\gamma d^{\prime}n^{\prime}}$. Recall that $n^{\prime\prime}\geq(1-2p)n^{\prime}\geq n^{\prime}/2$ and $d^{\prime}\geq\varepsilon d$, so for $\gamma^{\prime}:=\sqrt{96\gamma\varepsilon^{-2}K}$, we have
    $|S|\leq\sqrt{3\gamma d^{\prime}n^{\prime}}\leq\sqrt{6\gamma d^{\prime}n^{\prime\prime}}\leq\gamma^{\prime}d^{\prime},$
    where in the final inequality we used $Kd\geq\varepsilon n^{\prime\prime}/16$.
    Also note that $\gamma\leq\gamma^{\prime}\ll\varepsilon$ since we chose $\gamma\ll\varepsilon,1/K\ll 1$. Deleting vertices of $S$ from $V(G^{\prime\prime})$, we obtain $G^{\prime\prime\prime}:=G^{\prime\prime}\setminus S$ with minimum degree at least $(1-5\gamma^{\prime})d^{\prime}$. Indeed, for any vertex $v\in V(G^{\prime\prime\prime})$, we have that
    $\deg_{G^{\prime}}(v)=\deg_{G}(v;C^{\prime})-\deg_{G}(v,U^{c};C^{\prime})\geq d^{\prime}-\sqrt{3\gamma d^{\prime}n^{\prime}}\geq(1-\gamma^{\prime})d^{\prime}$
    since $v\notin S$, so by (9), we are sure to have $\deg_{G^{\prime\prime}}(v)\geq(1-3\gamma)(1-\gamma^{\prime})d^{\prime}\geq(1-4\gamma^{\prime})d^{\prime}$ since $\gamma\leq\gamma^{\prime}$. Finally, $v$ can send at most $|S|\leq\gamma^{\prime}d^{\prime}$ edges to $S$, so we are left with $\deg_{G^{\prime\prime\prime}}(v)\geq(1-5\gamma^{\prime})d^{\prime}$.
    Now $\delta(G^{\prime\prime\prime})\geq(1-5\gamma^{\prime})d^{\prime}\geq 9(\varepsilon/4)^{-3}$ since $d^{\prime}\geq\varepsilon d$, $1/d\ll\varepsilon$, and $\gamma^{\prime}\ll 1$. Therefore, we can apply Lemma 3.2 to the symmetric digraph associated to $G^{\prime\prime\prime}$ with $\varepsilon/4$ in place of $\varepsilon$ to obtain a rainbow path forest $\mathcal{P}$ with at most $9(\varepsilon/4)^{-2}=144\varepsilon^{-2}$ components and at least $(1-\varepsilon/4)(1-5\gamma^{\prime})d^{\prime}\geq(1-\varepsilon/2)d^{\prime}$ edges since $\gamma^{\prime}\ll\varepsilon$, as required. This proves the claim. ∎
Fix a vertex $u$ from the endpoints of $\mathcal{P}$. By deleting all edges with colours in $C(Q_{u})$ and all vertices in $V(Q_{u})\setminus\{u\}$ from $\mathcal{P}$, we obtain a rainbow path forest $\mathcal{P}^{\prime}$ whose colours are disjoint from $C(P\cup Q_{u})$, with $V(\mathcal{P^{\prime}})\cap V(P\cup Q_{u})=\{u\}$. Note that each vertex or colour we delete from $\mathcal{P}$ creates at most one new component and deletes at most two edges. Therefore, since $Q_{u}$ has length at most $s_{0}$, $d^{\prime}\geq\varepsilon d$, and $1/d\ll 1/s_{0}\ll\varepsilon$, we see that $\mathcal{P}^{\prime}$ has at most $144\varepsilon^{-2}+|C(Q_{u})|+|V(Q_{u})|-1\leq 3s_{0}$ components and at least $(1-\varepsilon/2)d^{\prime}-2(|C(Q_{u})|+|V(Q_{u})|-1)\geq(1-\varepsilon)d^{\prime}$ edges.
So far, we have a rainbow path $P\cup Q_{u}$ of length at least $\ell$ and a rainbow path forest $\mathcal{P}^{\prime}$ with at least $(1-\varepsilon)d^{\prime}$ edges which are colour-disjoint and intersect only at the shared endpoint $u$. We now join the components of $\mathcal{P}^{\prime}$ via short, internal-vertex-disjoint, colour-disjoint, rainbow paths using vertices from $V_{0}$ and colours from $C_{0}$ to obtain a single rainbow path of length at least $\ell+(1-\varepsilon)d^{\prime}\geq(1-\varepsilon)d$. Note that the only vertices and colours of $P\cup Q_{u}\cup\mathcal{P}^{\prime}$ that might appear in $V_{0}$ or $C_{0}$ come from the short path $Q_{u}$, so avoiding these when finding these connecting paths is not too difficult by B3. We denote the paths of $\mathcal{P^{\prime}}$ to be $P_{1},P_{2},\dots,P_{m}$, where $m\leq 3s_{0}$, and $P_{i}$ joins vertices $u_{i}$ and $v_{i}$, with $u_{1}=u$. Denote $Q_{0}:=Q_{u}$. For each $i=1,\ldots,m-1$, in turn, we iteratively apply property B3 of Lemma 4.3 to $v_{i}$ and $u_{i+1}$, with $V_{i}:=V_{0}\cap\cup_{j\leq i-1}V(Q_{j})$ and $C_{i}:=C_{0}\cap\cup_{j\leq i-1}C(Q_{j})$ playing the respective roles of $V_{1}$ and $C_{1}$, to obtain a rainbow path $Q_{i}$ connecting $v_{i}$ and $u_{i+1}$ of length at most $\nu^{-1}+1$ whose internal vertices are in $V_{0}\setminus V_{i}$ and whose colours are in $C_{0}\setminus C_{i}$. Note that at every step, we have $|V_{i}|,|C_{i}|\leq s_{0}+(\nu^{-1}+1)(m-1)\leq\beta n^{\prime}$, since $\beta=\gamma^{3}\nu/100$, $n^{\prime}\geq\delta(G^{\prime})\geq(1-3\gamma)d^{\prime}/2\geq\varepsilon d/4$, and $1/d\ll\nu,s_{0},\gamma,\varepsilon$, so the iterative application is valid. At the end we obtain a rainbow path $P\cup Q_{u}\cup\mathcal{P}^{\prime}\cup\ \bigcup_{i=1}^{m-1}Q_{i}$ of length at least $(1-\varepsilon)d$. ∎
7 The rearrangement problem in general groups
In this section, given a Cayley graph $G=\mathrm{Cay}(\Gamma,S)$ with $|S|=d$ sufficiently large, we show how to build a directed rainbow walk of length $d$ with at most $\varepsilon d$ vertex repetitions, thereby proving Theorem 1.3 (restated as Theorem 7.1 below). We achieve this incrementally, in $O(\varepsilon^{-1})$ steps. At a given stage $i$ in our algorithm, we have a directed rainbow walk $P_{i-1}$ and a set $S_{i-1}$ of unused colours. If the graph induced by the colour set $S_{i-1}$ has good expansion properties, then we can reach a large set $X_{i}$ of vertices by short (length $O(\varepsilon^{-1})$) $S_{i-1}$-rainbow paths from the terminal vertex of $P_{i-1}$, similar to the proof of Theorem 1.5 in Section 6. We can further ensure that all of these short paths use a common colour set $C_{i}$ of size $o(d)$. By Observation 2.3, we can find a rainbow path $Q_{i}$ of length $\Omega(\varepsilon d)$ in the remaining colours (possibly reusing vertices of $P_{i-1}$). By a simple double counting argument, there exists a translate of $Q_{i}$ which starts at a vertex in $X_{i}$ and intersects $P_{i-1}$ in $o(|Q|)$ vertices. We can thus append $Q_{i}$ to $P_{i-1}$ to obtain the new walk $P_{i}$, and we repeat. If, at some stage, we don’t have the necessary expansion properties to continue, then we can find a set $U$ of vertices in which $G[U;S_{i-1}]$ satisfies the hypotheses of Theorem 5.2. At this point, we can find a $S_{i-1}$-rainbow path $Q$ of length at least $(1-o(1))|S_{i-1}|$, which we can similarly translate to begin in the set $X_{i-1}$ from the previous iteration so that it only intersects $P_{i-2}$ in $o(|Q|)$ vertices. We also take care to ensure that $Q$ does not use colours from the set $C_{i-1}$ reserved for joining $P_{i-2}$ to $X_{i-1}$, so we can append $Q$ to $P_{i-2}$ to obtain a rainbow walk of length at least $(1-\varepsilon/2)d$ with at most $\varepsilon d/2$ vertex repetitions. Now an arbitrary extension in the remaining $\leq\varepsilon d/2$ colours gives the desired rainbow walk of length $d$.
Theorem 7.1.
For every $\varepsilon>0$, there exists $d_{0}\in\mathbb{N}$ such that for all groups $\Gamma$ and all subsets $S\subseteq\Gamma$ with $|S|=d\geq d_{0}$, $\mathrm{Cay}(\Gamma,S)$ contains a directed rainbow walk spanning $d$ edges with at most $\varepsilon d$ vertex repetitions.
  • Proof.
We choose constants $\gamma^{\prime},\gamma,\alpha,\theta,\delta$ and $\ell,k$ so that $1/n\leq 1/d\ll 1/\ell\ll\gamma^{\prime}\ll\gamma\ll\alpha\ll\theta,1/k,\delta\ll\varepsilon$. Let $S$ be a subset of size $d$ of a group $\Gamma$ of order $n$, and let $G:=\mathrm{Cay}(\Gamma,S)$. Since we are aiming for an asymptotic result, we may assume that $S$ does not contain the identity element.
Before we begin, we note that it suffices to find a rainbow directed walk $P$ of length at least $(1-\varepsilon/2)d$ with at most $\varepsilon d/2$ vertex repetitions. Then because every vertex has an out-neighbour in every colour from $S$, we can greedily extend $P$ to a rainbow walk of length $d$ with at most $\varepsilon d$ vertex repetitions.
We attempt to iteratively create rainbow walks $P_{0}\subseteq P_{1}\subseteq\cdots$ such that each $P_{i-1}$ is the initial segment of $P_{i}$. We also ensure that for each $i\geq 0$, we have $\delta d\leq|C(P_{i})|-|C(P_{i-1})|\leq\lceil\delta d\rceil+\ell$ and $|V(P_{i})|\geq|C(P_{i})|-2i\delta d/k-i\ell+1$. We terminate this process once we have $|C(P_{i})|\geq(1-\varepsilon/2)d$, unless our procedure fails before that point. Thus we will always have $i\leq\delta^{-1}$. At step $i=0$, let $v_{0}$ be an arbitrary vertex of $V(G)$ and $P_{0}$ be an empty walk with a single vertex $v_{0}$. For all $i$, $v_{0}$ will be the initial vertex of $P_{i}$, and we will denote the terminal vertex of $P_{i}$ by $v_{i}$. Suppose we are at step $i$ and that we have constructed $P_{0}\subseteq\cdots\subseteq P_{i-1}$. If we have that $|C(P_{i-1})|\geq(1-\varepsilon/2)d$, then we are already done with $P=P_{i-1}$, as the number of vertex repetitions of $P_{i-1}$ is only
$|C(P_{i-1})|+1-|V(P_{i-1})|\leq 2d/k+\delta^{-1}\ell\leq\varepsilon d/2,$
since $1/d\ll 1/\ell,1/k,\delta\ll\varepsilon$. Otherwise, we have $|S\setminus C(P_{i-1})|\geq\varepsilon d/2$, and we attempt to construct $P_{i}$ as follows.
We say that a colour set $C\subseteq S\setminus C(P_{i-1})$ is $(\theta,k,\ell)$-expanding if $|C|\leq\theta d$, and there exist at least $kd$ distinct vertices which are the terminal vertex of some $C$-rainbow walk of length at most $\ell$ from $v_{i-1}$.
If there exists a $(\theta,k,\ell)$-expanding colour set $C_{i}\subseteq S\setminus C(P_{i-1})$, let $X_{i}$ be the set of at least $kd$ vertices guaranteed by the definition of $(\theta,k,\ell)$-expanding. Because $|S\setminus(C(P_{i-1})\cup C_{i})|\geq\varepsilon d/2-\theta d\geq 2\lceil\delta d\rceil$ since $\delta,\theta\ll\varepsilon$, we can, by Observation 2.2, find a rainbow path $Q_{i}$ of length $\lceil\delta d\rceil$ starting at the identity element $e$ using colours from $S\setminus(C(P_{i-1})\cup C_{i})$. Next we will show that we can translate this path so that it starts at some vertex in $X_{i}$ without intersecting $P_{i-1}$ too much.
Claim.
There exists $x_{i}\in X_{i}$ such that $|x_{i}V(Q_{i})\cap V(P_{i-1})|\leq 2\delta d/k$.
  • Proof.
    We count the number $s$ of triples $(x,u,v)\in X_{i}\times V(Q_{i})\times V(P_{i-1})$ such that $xu=v$ in $\Gamma$. Clearly
    $s\leq|V(Q_{i})||V(P_{i-1})|\leq 2\delta d^{2}$
    since $|V(Q_{i})|\leq 1+\lceil\delta d\rceil\leq 2\delta d$ and $|V(P_{i-1})|\leq d$. On the other hand, if the claim does not hold, then
    $s>(2\delta d/k)|X_{i}|\geq 2\delta d^{2}$
    since $|X_{i}|\geq kd$, so we have a contradiction. ∎
Let $x_{i}$ be as in the claim, and let $L_{i}$ be a $C_{i}$-rainbow walk of length at most $\ell$ from $v_{i-1}$ to $x_{i}$. We take $P_{i}$ to be $P_{i-1}\cup L_{i}\cup(x_{i}Q_{i})$. By design, $P_{i-1}$, $L_{i}$, and $x_{i}Q_{i}$ are colour-disjoint, so $P_{i}$ is a rainbow walk with initial segment $P_{i-1}$. Moreover, $|C(P_{i})|-|C(P_{i-1})|=|C(L_{i})|+|C(Q_{i})|$ is between $\delta d$ and $\lceil\delta d\rceil+\ell$, and
$|V(P_{i})|$ $\geq|V(P_{i-1})|+|V(Q_{i})|-|x_{i}V(Q_{i})\cap V(P_{i-1})|$
$\geq|C(P_{i-1})|-2(i-1)\delta d/k-(i-1)\ell+1+|C(Q_{i})|-2\delta d/k$
$=|C(P_{i})|-|C(L_{i})|-2(i-1)\delta d/k-(i-1)\ell+1-2\delta d/k$
$\geq|C(P_{i})|-2i\delta d/k-i\ell+1,$
as required.
Thus we can assume that at some step $i$, our process fails because there exists no $(\theta,k,\ell)$-expanding set $C\subseteq S\setminus C(P_{i-1})$. We will use this fact to find an induced subgraph of our digraph which is dense in a large colour subset of
$S^{\prime}:=\left\{\begin{array}[]{l l}S&\text{if $i=1$,}\\ S\setminus(C(P_{i-1})\cup C_{i-1})&\text{if $i\geq 2$}.\end{array}\right.$
Recall that $v_{i-1}$ is the terminal vertex of $P_{i-1}$ and that
$|S^{\prime}|\geq|S\setminus C(P_{i-1})|-\theta d\geq\varepsilon d/2-\theta d\geq\theta d/2$
since $\theta\ll\varepsilon$. So there exists a set $Z$ consisting of $\lceil\theta d/2\rceil$ out-neighbours $w$ of $v_{i-1}$ for which $C(v_{i-1},w)\in S^{\prime}$. Let $C^{\prime}$ be the set of colours used on these edges. If there exists a colour $c\in S\setminus(C(P_{i-1})\cup C^{\prime})$ with at least $\gamma^{\prime}d$ $c$-edges coming out of $Z$, add the terminal vertices of those edges to $Z$ and add $c$ to $C^{\prime}$. Repeat this process as long as possible. Note that the size of $Z$ cannot ever exceed $kd$. If it did, then this would first occur after at most $kd/(\gamma^{\prime}d)+1\ll\ell$ iterations, so we would have $|C^{\prime}|\leq\lceil\theta d/2\rceil+\ell\leq\theta d$ since $1/d\ll 1/\ell,\theta$. Also every vertex in $Z$ would be reachable via a $C^{\prime}$-rainbow walk of length at most $\ell$ from $v_{i-1}$. Therefore, $C^{\prime}$ would be $(\theta,k,\ell)$-expanding, a contradiction. Thus, when this process terminates, we have that $\theta d/2\leq|Z|\leq kd$, that $|C^{\prime}|\leq\lceil\theta d/2\rceil+kd/(\gamma^{\prime}d)\leq\theta d$ since $1/d\ll\gamma^{\prime},\theta,1/k$, and that every colour in $C^{\prime\prime}:=S^{\prime}\setminus C^{\prime}$ has at most $\gamma^{\prime}d$ edges going from $Z$ to $Z^{c}$. The size of $C^{\prime\prime}$ is at least $|C^{\prime\prime}|\geq d-\theta d$ if $i=1$ since $|C^{\prime}|\leq\theta d$. If $i\geq 2$, then we have $|C^{\prime\prime}|\geq\varepsilon d/2-2\theta d\geq\varepsilon d/4$ since $|C_{i-1}|\leq\theta d$, and $|S\setminus C(P_{i-1})|\geq\varepsilon d/2$, and since $\theta\ll\varepsilon$. From this, we can deduce a stronger lower bound on $|Z|$ than $\theta d/2$; because $\gamma^{\prime}d\leq\theta d/4\leq|Z|/2$, we have
$\varepsilon d|Z|/8\leq|C^{\prime\prime}|(|Z|-\gamma^{\prime}d)\leq e(G[Z;C^{\prime\prime}])\leq|Z|^{2},$
and therefore $|Z|\geq\varepsilon d/8$.
Define $r:=\sqrt{\gamma^{\prime}}d$. Recall that for each $c\in C^{\prime\prime}$, the colour class of $c$ in $G$ is a $1$-factor with at most $\gamma^{\prime}d$ $c$-edges from $Z$ to $Z^{c}$, so there are also at most $\gamma^{\prime}d$ $c$-edges from $Z^{c}$ to $Z$. Consequently, the set $Y$ of vertices in $Z$ with in-degree or out-degree below $|C^{\prime\prime}|-r$ in $G[Z;C^{\prime\prime}]$ has size at most
$|Y|\leq\frac{2\gamma^{\prime}d|C^{\prime\prime}|}{r}=2\sqrt{\gamma^{\prime}}|C^{\prime\prime}|\leq 2r.$
Now $G^{\prime}:=G[Z\setminus Y;C^{\prime\prime}]$ satisfies the following properties. Let $n^{\prime}:=|V(G^{\prime})|$. First,
$kd\geq n^{\prime}=|Z|-|Y|\geq\left(\varepsilon/8-2\sqrt{\gamma^{\prime}}\right)d\geq\varepsilon d/16$
since $\gamma^{\prime}\ll\varepsilon$. Consequently,
$\delta^{\pm}(G^{\prime})\geq|C^{\prime\prime}|-3r\geq|C^{\prime\prime}|-48\left(\sqrt{\gamma^{\prime}}/\varepsilon\right)n^{\prime}\geq|C^{\prime\prime}|-\gamma n^{\prime}\geq\Delta^{\pm}(G^{\prime})-\gamma n^{\prime}$
since $\gamma^{\prime}\ll\gamma,\varepsilon$. Also, the size of $C^{\prime\prime}$ is at least
$|C^{\prime\prime}|\geq\varepsilon d/4\geq\varepsilon n^{\prime}/(4k)\geq\alpha n^{\prime}+\gamma n^{\prime}$
since $\gamma,\alpha\ll 1/k,\varepsilon$, and so $\delta^{\pm}(G^{\prime})\geq\alpha n^{\prime}$.
Since $\gamma\ll\alpha,\varepsilon$, we can apply Theorem 5.2 with $\varepsilon/4$, $n^{\prime}$, $G^{\prime}$ in place of $\varepsilon$, $n$, $G$, respectively, to obtain a $C^{\prime\prime}$-rainbow directed path $Q$ of length at least $(1-\varepsilon/4)\delta^{\pm}(G^{\prime})\geq(1-\varepsilon/4)(|C^{\prime\prime}|-\gamma n^{\prime})$ in $G^{\prime}$. If $i=1$, then we can take $P=Q$, which is already a rainbow walk of length at least
$(1-\varepsilon/4)((1-\theta)d-\gamma n^{\prime})\geq(1-\varepsilon/4)(1-\theta-\gamma k)d\geq(1-\varepsilon/2)d$
with no vertex repetitions, as $\theta\ll\varepsilon$, and $\gamma\ll 1/k,\varepsilon$.
Otherwise we have $i\geq 2$, and we can append a translate of $Q$ to the walk $P_{i-2}$ found previously. Let $w$ be the initial vertex of $Q$. Recall that by construction, there exists a set $X_{i-1}$ of at least vertices such that each vertex $x\in X_{i-1}$ is reachable from $v_{i-2}$ by a $C_{i-1}$-rainbow path of length at most $\ell$.
Claim.
There exists $x\in X_{i-1}$ such that $|xw^{-1}V(Q)\cap V(P_{i-2})|\leq\varepsilon d/4$.
  • Proof.
    Similar to the previous claim, we count the number $s$ of triples $(x,u,v)\in X_{i-1}\times V(Q)\times V(P_{i-2})$ such that $xw^{-1}u=v$ in $\Gamma$. Recall that $Q$ and $P_{i-2}$ are both nonempty, rainbow, and colour-disjoint, so each has at most $d-1$ edges, hence at most $d$ vertices. Thus
    $s\leq|V(Q)||V(P_{i-2})|\leq d^{2}.$
    On the other hand, if the claim does not hold, then
    $s>(\varepsilon d/4)|X_{i-1}|\geq\varepsilon kd^{2}/4\geq d^{2}$
    since $|X_{i-1}|\geq kd$ and $1/k\ll\varepsilon$, a contradiction. ∎
Now, we let our walk be $P:=P_{i-2}\cup L\cup(xw^{-1}Q)$, where $L$ is a $C_{i-1}$-rainbow path of length at most $\ell$ joining $v_{i-2}$ to $x$. Note that $C(P_{i-2})$, $C_{i-1}$, and $C^{\prime\prime}$ are disjoint colour sets by design, so $P$ is indeed a rainbow walk. Also note that the length of $P$ is at least
$|C(P_{i-2})|+(1-\varepsilon/4)(|C^{\prime\prime}|-\gamma n^{\prime})$
$\geq|C(P_{i-1})|-\ell-\lceil\delta d\rceil+(1-\varepsilon/4)(d-|C(P_{i-1})|-2\theta d-\gamma kd)$
$\geq(1-\varepsilon/2)d$
since $1/d\ll 1/\ell,\theta,\delta\ll\varepsilon$, and $\gamma\ll 1/k,\varepsilon$. Finally, $P$ has at most
$|C(P)|-|V(P)|+1$ $\leq|C(P)|-|V(P_{i-2})|-|V(Q)|+|xw^{-1}V(Q)\cap V(P_{i-2})|+1$
$\leq|C(P_{i-2})|-|V(P_{i-2})|+1+|C(L)|+|C(Q)|-|V(Q)|+\varepsilon d/4$
$\leq 2(i-2)\delta d/k+(i-2)\ell+\ell+\varepsilon d/4$
$\leq 2d/k+\delta^{-1}\ell+\varepsilon d/4\leq\varepsilon d/2$
vertex repetitions since $1/d\ll 1/\ell,1/k,\delta\ll\varepsilon$. ∎
Remark. The proof above uses only basic properties of groups and can be easily adapted to any vertex-transitive digraph with a vertex-transitive proper edge colouring; i.e., properly edge-coloured digraphs with the property that for any two vertices $u$ and $v$, there is a colour-preserving automorphism of the digraph mapping $u$ to $v$.
8 Concluding remarks
Rearrangeable subsets in groups. In Theorem 1.4(b), we showed that dense coloured Cayley graphs contain directed rainbow paths with a length that is asymptotically best possible. Proving a similar result without a density assumption would be of great interest.
Problem 8.0.
For any group $\Gamma$ of order $n$ and any subset $S\subseteq\Gamma$ of size $d$, show that there exists a subset $S^{\prime}\subseteq S$ of size $d-o(d)$ which is rearrangeable.
Problem 8 is already open for cyclic groups $\mathbb{Z}_{p}$ of prime order. In that case, Theorem 1.4(c) finds the desired subset when $d\geq p^{3/4+o(1)}$. By results of [11], the entire set $S$ is rearrangeable as long as $d\leq e^{(\log p)^{1/4}}$ and $0\notin S$. We know how to improve the first result to lower density, namely $d\geq p^{2/3+o(1)}$, with a more involved argument which we chose to omit in this paper for the sake of brevity. However, our methods meet a natural barrier at roughly $d\sim p^{1/2}$, as then the diameter of $\mathrm{Cay}(\mathbb{Z}_{p},S)$ can be larger than the total number of colours available, so a lemma analogous to Lemma 4.3 cannot hold.
Rainbow walks in regular digraphs. Regarding general coloured $d$-regular digraphs, our current methods appear to be too weak to give a positive answer to Section 1. We pose the following relaxation as a more approachable open problem, with Theorem 1.3 already giving a positive answer for coloured Cayley graphs.
Problem 8.0.
Let $G$ be a $d$-regular digraph properly edge-coloured with $d$ colours. Does $G$ contain a rainbow walk with $d-o(d)$ distinct vertices?
Rainbow paths of length asymptotic in minimum degree. On the other hand, in proving the asymptotic results Theorem 6.4 and Theorem 5.2, we were able to drop the condition on the number of colours and relax the regularity condition to some extent. We wonder if a minimum out-degree assumption is enough to guarantee similar results.
Problem 8.0.
Let $G$ be properly edge-coloured digraph with minimum out-degree $d$. Does $G$ contain a rainbow path of length $d-o(d)$?
Note that Observation 2.2 gives a rainbow path of length at least $d/2$ under these conditions, and we know of no improvement to this simple bound for directed graphs. In the undirected setting, Johnston, Palmer, and Sarkar [38] showed that any properly edge-coloured graph of minimum degree $d$ contains a rainbow path of length at least $32d/$ 3. Chen and Li [17] considered the more general setting of graphs with arbitrary edge colourings in which every vertex is incident to edges with at least $d$ distinct colours. They proved that in such a regime, it is always possible to find a rainbow path of length at least $12d/3+$ 1 in general (see [8]), and at least $d-1$ if $d\leq 7$. They went on to conjecture that a rainbow path of length $d-1$ always exists. Even proving a bound of $d-o(d)$ in either of these undirected settings would be interesting.
Rainbow Turán numbers of paths. We may also consider analogous problems for average degree. Specifically, given $d>0$, what is the largest integer $f(d)$ such that every properly edge-coloured graph with average degree at least $d$ contains a rainbow path of length $f(d)$? This is equivalent to the so-called rainbow Turán problem for paths, formally introduced in the influential paper by Keevash, Mubayi, Sudakov and Verstraëte [42]. In the uncoloured setting, Erdős and Gallai [23] showed that any graph with average degree $d$ contains a path of length at least $d$, and this is best possible when $d$ is an integer by considering a disjoint union of cliques of size $d+1$. In comparison, no tight bound is known for $f(d)$ despite considerable effort in this direction [36, 35, 38, 25, 39]. Note that Observation 2.3 gives an easy bound of $f(d)\geq d/4$. The current best known lower bound is $f(d)\geq 7d/18-O(1)$ by Ergemlidze, Györi and Methuku [25], and the best upper bound is $f(d)\leq\lceil d\rceil-1$ for $d>2$ by Johnston and Rombach [39], coming from the coloured Cayley graph $\mathrm{Cay}(\mathbb{F}_{2}^{k},S)$ generated by a zero-sum set $S$ of size $\lceil d\rceil$. This upper bound is known to give the correct value of $f(d)$ for $2<d\leq 6$ [38, 35], and Halfpap [35] has conjectured it to be tight for all $d$, up to possibly an additive constant.
The natural analogue of this question for general directed graphs is somewhat less interesting, as directed bicliques can have very high average out-degree but no directed path of length $2$, let alone a rainbow one. To remedy this, one may impose additional assumptions on the digraph. One natural condition to consider is that the digraph is Eulerian, meaning that every vertex has in-degree equal to its out-degree.
Problem 8.0.
What is the maximum length of a rainbow path in any properly edge-coloured Eulerian digraph of average out-degree $d$?
In the uncoloured setting, a long-standing conjecture of Bollobás and Scott (Conjecture 7 in [14]) implies that any Eulerian digraph of average out-degree $d$ contains a path of length $\Omega(d)$. The best known lower bound in this direction is of order $\Omega(d/\log{d})$ by Knierim, Larcher and Martinsson [45]. Our Lemma 4.2 might provide some insight into questions such as this, as it gives a way to find robust out-expanders in dense Eulerian digraphs. However, it seems to say nothing without any minimum degree assumptions, so additional ideas are needed.
Acknowledgments. We would like to thank Noga Alon for leading us to Pollard’s Theorem which replaced a more involved earlier argument for establishing expansion properties in $\mathbb{Z}_{p}$.